### v1

# Quadratic dispersion relations in gapless frustration-free systems

Abstract
Recent case-by-case studies revealed that the dispersion of low energy excitations in gapless frustration-free Hamiltonians is often quadratic or softer. In this work, we argue that this is actually a general property of such systems. By combining a previous study by Bravyi and Gosset and the min-max principle, we prove this hypothesis for models with local Hilbert spaces of dimension two that contains only nearest-neighbor interactions on cubic lattice. This may be understood as a no-go theorem realizing gapless phases with linearly dispersive excitations in frustration-free Hamiltonians. We also provide examples of frustration-free Hamiltonians in which the plane-wave state of a single spin flip does not constitute low energy excitations.
I Introduction
Accurately investigating the ground states and excited states of quantum many-body systems is often challenging. It is hence common practice to analyze the system of interest using the simplest possible model that belongs to the same phase. Frustration-free systems, in particular, are relatively easy to handle, and their exact ground states and excited states are sometimes accessible. For these reasons, frustration-free systems have been widely used as representative models of various phases of matter. Therefore, elucidating the general properties of frustration-free systems and understanding the limits of the phases that can be realized by frustration-free Hamiltonians hold significant importance.
Numerous frustration-free systems posses an excitation gap. Notable examples include the Majumdar–Ghosh (MG) model, known as a toy model realizing spontaneous breaking of discrete translational symmetry, the Affleck–Kennedy–Lieb–Tasaki (AKLT) model, representing a symmetry-protected topological phase, and the Kitaev Toric code model, serving as the canonical model for topologically ordered phases Auerbach (2012); Tasaki (2020). The parent Hamiltonians of Matrix Product States (MPS) also feature an excitation gap Fannes et al. (1992).
There are also frustration-free Hamiltonians with gapless excitations. For example, the uncle Hamiltonian of MPS Fernández-González et al. (2015), various models at critical points, and models related to spontaneously broken continuous symmetries Ogunnaike et al. (2023); Watanabe et al. (2023) are gapless. Through recent investigation of these models, the general properties of frustration-free systems have gradually become clearer.
In this paper, we discuss two conjectures regarding frustration-free systems. The first is about the finite-size gap between degenerate ground states. Generally, even if the ground states are degenerate in the thermodynamic limit, their energy eigenvalues in a finite-size system do not precisely match due to the finite-size gap. However, in frustration-free systems, the ground states are exactly degenerate even before taking the thermodynamic limit. In other words, we can judge whether a frusration-free system is gapped or not by looking at the limiting behavior of the difference between the ground state energy and the second lowest energy level.
The second conjecture concerns the dispersion relation of low-energy excitations in frustration-free systems. Although gapless modes typically have a linear dispersion, it is known that the ferromagnetic Heisenberg model, which is a representative frustration-free Hamiltonian, have a spin wave excitation with a quadratic dispersion Auerbach (2012). In fact, our second conjecture asserts that gapless frustration-free systems generally have a quadratic or softer dispersion relation. Although there are several arguments supporting this statement in literature (for example, see Refs. Palle and Benton, 2021; Ogunnaike et al., 2023; Watanabe et al., 2023; Ren et al., 2024), they are applicable only to the case where the low-energy excitations are obtained by applying a sum of local operators to a ground state. We provide a general proof for spin-1/2 models with nearest-neighbor interaction on cubic lattice based on the result of Bravyi and Gosset Bravyi and Gosset (2015). We also discuss examples with longer-range interactions and show that gapless excitations with quadratic dispersion may not be obtained by a sum of local operators.
II Overview
II.0.1 Setting and definitions
In this work, we consider Hamiltonians defined on a finite $d$-dimensional lattice $\Lambda$:
$\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{H}_{\bm{r}}.$ (1)
The local Hilbert space on each lattice site $\bm{r}\in\Lambda$ is assumed to be finite dimensional, which is denoted by $D_{0}$. Let $|\Lambda|$ be the number of lattice sites in $\Lambda$. Then the dimension of the total Hilbert space is $D=D_{0}^{|\Lambda|}$. We assume that the Hamiltonian is local in the sense that $\hat{H}_{\bm{r}}$ acts nontrivially on the sites within a finite distance $R$ from $\bm{r}$.
Let us write eigenvalues of $\hat{H}$ as
$E_{1}\leq E_{2}\leq\cdots\leq E_{D}.$ (2)
The thermodynamic limit $|\Lambda|\to\infty$ is taken by a sequence of increasing system size.
The system is said to be gapped if there exists an integer $N_{\mathrm{deg}}$ ($1\leq N_{\mathrm{deg}}<D$), which may depend on the system size, such that
$\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}}-E_{1}\big{)}=0,$ (3)
$\Delta\coloneqq\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}+1}-E_{N_{\mathrm{deg}}}\big{)}\neq 0.$ (4)
Otherwise the system is gapless. The limit in Eq. (3) defines the excitation gap $\Delta$. When $N_{\mathrm{deg}}\geq 2$, the quantity $E_{N_{\mathrm{deg}}}-E_{1}$ is referred to as the finite-size gap among degenerate ground states. When the degeneracy originates from spontaneous breaking of a discrete symmetry or formation of a topological order, it generically decays exponentially with the system size. Even when $N_{\mathrm{deg}}$ converges to a finite number in the large $|\Lambda|$ limit, its value for a finite system may be ambiguous due to the possible system-size dependence.
The Hamiltonian is translation invariant if there exists an operator $\hat{T}_{\bm{a}}$ for each lattice vector $\bm{a}$ such that $\hat{T}_{\bm{a}}\hat{H}_{\bm{r}}=\hat{H}_{\bm{r}+\bm{a}}\hat{T}_{\bm{a}}$ for all $\bm{r}\in\Lambda$. This is possible only when the periodic boundary condition (PBC) is imposed.
The Hamiltonian is frustration-free if a ground state $|{\Phi_{0}}\rangle$ of $\hat{H}$ is a simultaneous ground state of every local Hamiltonian $\hat{H}_{\bm{r}}$ Tasaki (2020). In this case, without loss of generality and without modifying the ground states and low-energy excitations of the system, one may assume that each $\hat{H}_{\bm{r}}$ is a projector, i.e., $\hat{Q}_{\bm{r}}^{2}=\hat{Q}_{\bm{r}}$ and $\hat{Q}_{\bm{r}}|{\Phi_{0}}\rangle=0$ for all $\bm{r}\in\Lambda$. Under this choice, the ground state energy becomes $E_{1}=0$.
II.0.2 Conjectures on frustration-free Hamiltonian
Let us state our conjectures one by one. The first conjecture is that, when the Hamiltonian is frustration-free and gapped, the finite-size gap among ground states precisely vanishes before taking the thermodynamic limit:
$E_{N_{\mathrm{deg}}}=E_{1}.$ (5)
For example, in the AKLT model for spin $s=1$ chain under OBC, the four-fold ground state degeneracy due to the edge modes is exact. The four-fold topological degeneracy in the Kitaev toric code under PBC is also exact.
This implies the following criterion for the excitation gap in frustration-free systems. Let $\tilde{N}_{\mathrm{deg}}$ ($1\leq\tilde{N}_{\mathrm{deg}}<D$) be the number of exact zero energy states, i.e.,
$E_{\tilde{N}_{\mathrm{deg}}}=E_{1},\quad E_{\tilde{N}_{\mathrm{deg}}+1}\neq E_{1},$ (6)
which may also depend on the system size (see, for example, Eq. (45) below). The Hamiltonian is gapped if and only if
$\lim_{|\Lambda|\to\infty}\big{(}E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\big{)}>0.$ (7)
Namely, one can set $N_{\mathrm{deg}}=\tilde{N}_{\mathrm{deg}}$. Actually this criterion has been used as the definition of the excitation gap in previous studies (for example, see Refs. Bravyi and Gosset, 2015; Gosset and Mozgunov, 2016; Gosset and Huang, 2016). Note that $\tilde{N}_{\mathrm{deg}}$ is well-defined even for gapless systems unlike $N_{\mathrm{deg}}$.
The second conjecture is that, when the Hamiltonian is translation-invariant, frustration-free, and gapless, there exists a family of variational states $|\Psi_{\bm{k}}\rangle$ that are orthogonal to all the ground states and satisfy
$\hat{T}_{\bm{a}}|\Psi_{\bm{k}}\rangle=e^{-i\bm{k}\cdot\bm{a}}|\Psi_{\bm{k}}\rangle,$ (8)
$\langle\Psi_{\bm{k}}|\hat{H}|\Psi_{\bm{k}}\rangle=O\big{(}|\bm{k}-\bm{k}_{0}|^{2},L^{-2}\big{)}$ (9)
for some $\bm{k}_{0}$. Namely, the low energy excitation of the system is not linearly dispersive about the gapless point at $\bm{k}=\bm{k}_{0}$.
For systems under OBC, the result of Refs. Knabe, 1988; Anshu, 2020; Gosset and Mozgunov, 2016 gives us the bound $E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\leq O(L^{-2})$. Although this is consistent with our conjecture, it is not directly applicable to the PBC case Watanabe et al. (2023).
II.0.3 Min-max principle
Next, let us review the min-max principle, which supports our conjectures. We compare two Hamiltonians $\hat{H}$ and $\hat{H}^{\prime}$ acting on a $D$-dimensional Hilbert space. The eigenvalues of $\hat{H}$ and $\hat{H}^{\prime}$ are, respectively, denoted by $E_{j}$ and $E_{j}^{\prime}$ ($j=1,2,\cdots,D$) in the increasing order. Suppose that $\hat{V}\coloneqq\hat{H}^{\prime}-\hat{H}$ is positive semi-definite:
$\hat{V}\geq 0.$ (10)
That is, $\langle\Phi|\hat{V}|\Phi\rangle\geq 0$ for any state $|\Phi\rangle$. In other words, all the eigenvalues of $\hat{V}$ are nonnegative. In this setting, the min-max principle states that
$E_{j}^{\prime}\geq E_{j}$ (11)
for every $j=1,2,\cdots,D$ (see Theorem A.7 of Ref. Tasaki (2020)). This statement is trivial when the two Hamiltonians can be diagonalized by a common unitary operator, but it holds more generally. Since the proof is elementary, we review it in Appendix A.
Two corollaries follow immediately from the min-max principle. Let us consider two frustration free Hamiltonians $\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}$ and $\hat{H}^{\prime}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}^{\prime}$. Suppose that $\hat{V}=\hat{H}^{\prime}-\hat{H}$ is positive-semidefinite. Then the inequality in (11) implies that the number of zero-energy states of $\hat{H}$ cannot be smaller than that of $\hat{H}^{\prime}$:
$\tilde{N}_{\mathrm{deg}}\geq\tilde{N}_{\mathrm{deg}}^{\prime}.$ (12)
Furthermore,
$\lim_{|\Lambda|\to\infty}E_{j}^{\prime}=0\quad\Rightarrow\quad\lim_{|\Lambda|\to\infty}E_{j}=0.$ (13)
This relation suggests that if (i) $\hat{H}$ is gapped and (ii) $\tilde{N}_{\mathrm{deg}}$ is bounded by a system-size independent constant, then $\hat{H}^{\prime}$ is also gapped. This relation was previously used for interacting Kitaev chains in Refs. Katsura et al., 2015; Wouters et al., 2018.
III 1D Examples with nearest-neighbor interactions
[…]
$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-1}\hat{Q}_{x,x+1}.$ (14)
Here, $\hat{Q}_{x,x+1}=\hat{Q}_{x,x+1}^{2}$ is a projector nontrivially acting on the spins at $x$ and $x+1$. The interaction range of this Hamiltonian is 1. The number of $+1$ eigenvalues of a projector is called the rank. The Hamiltonian $\hat{H}^{\mathrm{pbc}}$ has the translation symmetry $\hat{T}$ which satisfies $\hat{T}\hat{Q}_{x-1,x}=\hat{Q}_{x,x+1}\hat{T}$.
The seminal work by Bravyi and Gosset Bravyi and Gosset (2015) showed that $\hat{H}^{\mathrm{obc}}$ can be gapless only when the projector $\hat{Q}_{x,x+1}$ is unitary equivalent to Eqs. (19), (20) for the rank 1 case and Eq. (38) for the rank 2 case. Then the min-max principle implies the same for $\hat{H}^{\mathrm{pbc}}$ at least when $\tilde{N}_{\mathrm{deg}}$ under OBC is finite, since $\hat{H}^{\mathrm{pbc}}-\hat{H}^{\mathrm{obc}}=\hat{Q}_{L,1}\geq 0$ is positive semi-definite. We will see that the fully polarized state
$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x=1}^{L}|0\rangle_{x}=|0\cdots 0\rangle$ (15)
is a common ground state of these Hamiltonians. Furthermore, the plane-wave state of a single spin flip
$|{\Psi_{k}}\rangle\coloneqq\frac{1}{\sqrt{L}}\sum_{x=1}^{L}e^{ikx}|x\rangle,$ (16)
$|x\rangle\coloneqq\hat{s}_{x}^{-}|{\Phi_{0}}\rangle=|\underbrace{0\cdots 0}_{x-1}1\underbrace{0\cdots 0}_{L-x}\rangle$ (17)
with $k\coloneqq 2\pi m/L$ ($m=1,2,\cdots L$) is a common variational state among these examples and has a quadratic dispersion relation. That is,
$E(k)$ $=\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\sum_{x=1}^{L}\big{\|}\hat{Q}_{x,x+1}|{\Psi_{k}}\rangle\big{\|}^{2}$
$=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik(x-x^{\prime})}\langle x^{\prime}|\hat{H}^{\mathrm{pbc}}|x\rangle=O\big{(}(k-k_{0})^{2}\big{)}$ (18)
for some $k_{0}$.
III.1 Rank 1 case
Let us start with the rank 1 case. We write
$\hat{Q}_{x,x+1}\coloneqq|\psi\rangle_{x,x+1}\langle\psi|_{x,x+1},$ (19)
where $|\psi\rangle_{x,x+1}$ is a normalized state for two spins at $x$ and $x+1$. Without loss of generality one can assume the form
$|\psi\rangle=(\alpha+i\beta)|01\rangle+(\alpha+i\gamma)|10\rangle+\delta|11\rangle,$ (20)
$\alpha,\beta,\gamma,\delta\in\mathbb{R},\quad 2\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}=1$ (21)
by performing a local unitary transformation (see Appendix B for the derivation) Bravyi and Gosset (2015). The model has U(1) spin rotation symmetry about the $z$ axis if and only if $\delta=0$.
Since the matrix elements are known, the Hamiltonian in Eq. (19) can be expressed in terms of spin operators $\hat{s}_{x}^{a}$ and $\hat{s}_{x+1}^{a}$ ($a=x,y,z$). For example, when $(\alpha,\beta,\gamma,\delta)=\pm(0,1/\sqrt{2},-1/\sqrt{2},0)$, $\hat{Q}_{x,x+1}$ becomes the projector onto the spin-singlet state $(|01\rangle-|10\rangle)/\sqrt{2}$:
$\hat{Q}_{x,x+1}$ $=\hat{\Pi}_{x,x+1}^{s=0},$ (22)
$\hat{\Pi}_{x,x+1}^{s=0}$ $\coloneqq\hat{\mathbbm{1}}-\frac{1}{2}(\hat{\bm{s}}_{x}+\hat{\bm{s}}_{x+1})^{2}=\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (23)
The corresponding Hamiltonian is the ferromagnetic Heisenberg model. Any state that is invariant under permutations of any two qubits (e.g., the fully-polarized state and the W state) is a ground state of this model. As a basis of such states, we can choose
$\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L),$ (24)
where $\hat{\mathcal{S}}$ is the symmetrization operator, uniformly averaging over the $\binom{L}{n}=\frac{L!}{(L-n)!n!}$ states. This model is known to be gapless due to the presence of a spin-wave excitation $E(k)=2\sin^{2}(k/2)$ under PBC.
To analyze more general cases, let us introduce a matrix
$m_{\psi}\coloneqq\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}=\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}.$ (25)
The model defined by Eqs. (14), (19), and (20) can be gapless only when $\det m_{\psi}=(\alpha-i\beta)(-\alpha+i\gamma)\neq 0$, i.e., $|\psi\rangle$ is entangled Bravyi and Gosset (2015). In this case, we introduce an invertible operator $\hat{M}$ composed of nonuniform powers of the matrix $m_{\psi}$:
$\hat{M}\coloneqq\mathbbm{1}\otimes\hat{m}_{\psi}\otimes\hat{m}_{\psi}^{2}\otimes\cdots\otimes\hat{m}_{\psi}^{L-1},$ (26)
$\hat{m}_{\psi}\coloneqq(\alpha-i\beta)|0\rangle\langle 0|-(\alpha-i\gamma)|1\rangle\langle 1|+\delta|0\rangle\langle 1|.$ (27)
This operator maps $|\psi\rangle$ to the singlet state:
$\hat{M}^{\dagger}|\psi\rangle_{x,x+1}=(\det{m_{\psi}}^{*})^{j}(|01\rangle_{x,x+1}-|10\rangle_{x,x+1}),$ (28)
It follows that $\hat{Q}_{x,x+1}$ for the state in Eq. (20) can be connected to $\hat{\Pi}_{x,x+1}^{s=0}$ as
$\hat{Q}_{x,x+1}=4|{\det{m}_{\psi}}|^{2}(\hat{M}\hat{\Pi}_{x,x+1}^{s=0}\hat{M}^{-1})^{\dagger}(\hat{M}\hat{\Pi}_{x,x+1}^{s=0}\hat{M}^{-1}).$ (29)
This type of non-unitary transformation is known as Witten’s conjugation Witten (1982); Wouters et al. (2021). As a consequence, ground states of $\hat{H}^{\mathrm{obc}}$ takes the form Bravyi and Gosset (2015)
$\hat{M}\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L).$ (30)
The Hamiltonian under PBC has an additional term $\hat{Q}_{L,1}$. Every ground state of $\hat{H}^{\mathrm{obc}}$ remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if $m_{\psi}^{L}$ is proportional to the identity matrix $\mathbbm{1}$. However, except when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$ discussed above, $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$ for some $L$. In such a case, the ground state of $\hat{H}^{\mathrm{pbc}}$ is restricted to the product state of the eigenstates of $\hat{m}_{\psi}$; that is, $|{\Phi_{0}}\rangle=|00\cdots 0\rangle$ and $|uu\cdots u\rangle$, where
$|u\rangle\coloneqq\frac{[2\alpha-i(\beta+\gamma)]|1\rangle-\delta|0\rangle}{\sqrt{4\alpha^{2}+(\beta+\gamma)^{2}+\delta^{2}}}.$ (31)
One can easily check that $\langle\psi|00\rangle=\langle\psi|uu\rangle=0$. Therefore, the exact ground state degeneracy of this model can be summarized as Bravyi and Gosset (2015)
$\tilde{N}_{\mathrm{deg}}=\begin{cases}L+1&\text{if }m_{\psi}^{L}\propto\mathbbm{1},\\ 2&\text{otherwise}.\\ \end{cases}$ (32)
The variational state $|{\Psi_{k}}\rangle$ in Eq. (16) with $k\neq 0$ is orthogonal to the ground states at least when $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$, since their translation eigenvalues are different. The energy expectation value in Eq. (18) can be computed as
$E(k)$ $=\sum_{x=1}^{L}\big{|}\langle\psi|_{x,x+1}|{\Psi_{k}}\rangle\big{|}^{2}=\big{|}\langle\psi|10\rangle+e^{ik}\langle\psi|01\rangle\big{|}^{2}$
$=\big{|}(\alpha-i\gamma)+(\alpha-i\beta)e^{ik}\big{|}^{2}$
$=(1-\delta^{2})+2(\alpha^{2}+\beta\gamma)\cos k+2\alpha(\beta-\gamma)\sin k$
$=(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}\cos(k-k_{0}).$ (33)
In going to the last line, we defined
$\cos k_{0}\coloneqq-\frac{\alpha^{2}+\gamma\beta}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}},$ (34)
$\sin k_{0}\coloneqq-\frac{\alpha(\beta-\gamma)}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}}.$ (35)
It follows that $E(k)$ can be bounded below as
$E(k)\geq(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}.$ (36)
The right-hand side vanishes when $\beta^{2}=\gamma^{2}$, implying the existence of gapless excitations when $\gamma=\pm\beta$. In this case $E(k)$ can be rewritten as
$E(k)=2(1-\delta^{2})\sin^{2}((k-k_{0})/2),$ (37)
which confirms our conjecture in the rank 1 case. When $\gamma=\beta$, $k_{0}$ is precisely $\pi$, but it takes more general value when $\gamma=-\beta$. This result is consistent with Ref. Bravyi and Gosset (2015) which mathematically proved that $\hat{H}^{\mathrm{obc}}$ is gapped if and only if $\beta^{2}\neq\gamma^{2}$.
III.2 Rank 2 case
Next let us discuss the rank 2 case. In this case the Hamiltonian in Eq. (14) can be gapless only when $\hat{Q}_{x,x+1}$ is unitary equivalent to the XY model with an external magnetic field:
$\hat{Q}_{x,x+1}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\frac{\zeta\hat{s}_{x}^{+}\hat{s}_{x+1}^{-}+\zeta^{*}\hat{s}_{x}^{-}\hat{s}_{x+1}^{+}}{1+|\zeta|^{2}}-\frac{|\zeta|^{2}\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}}{{1+|\zeta|^{2}}}.$ (38)
with $\zeta\in\mathbb{C}$ Bravyi and Gosset (2015). One may express this Hamiltonian as
$\mathrm{ker}\,\hat{Q}_{x,x+1}=\mathrm{span}\big{\{}|00\rangle,|10\rangle+\zeta|01\rangle\big{\}}.$ (39)
When $\zeta=1$, ground states of $\hat{Q}_{x,x+1}^{\zeta=1}$ are the fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of single spin flip $|\Psi_{0}\rangle$, regardless of the boundary condition. We find that $\hat{Q}_{x,x+1}$ with $\zeta\neq 1$ can be obtained from $\hat{Q}_{x,x+1}^{\zeta=1}$ by Witten’s conjugation as
$\hat{Q}_{x,x+1}=(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1})^{\dagger}\hat{C}_{x,x+1}(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1}),$ (40)
where
$\hat{M}\coloneqq\hat{\mathbbm{1}}\otimes\hat{m}_{\zeta}\otimes\hat{m}_{\zeta}^{2}\otimes\cdots\otimes\hat{m}_{\zeta}^{L-1},$ (41)
$\hat{m}_{\zeta}\coloneqq|0\rangle\langle 0|+\zeta|1\rangle\langle 1|,$ (42)
$\hat{C}\coloneqq\hat{\mathbbm{1}}+\frac{(1-|\zeta|^{2})^{2}}{2\mathrm{Re}\zeta(1+|\zeta|^{2})}(|10\rangle\langle 01|+|01\rangle\langle 10|).$ (43)
Hence, the ground states of $\hat{H}^{\mathrm{obc}}$ with $\zeta\neq 1$ are given by $\hat{M}|{\Phi_{0}}\rangle=|{\Phi_{0}}\rangle$ and $\hat{M}|\Psi_{0}\rangle\propto\sum_{x=1}^{L}\zeta^{x-1}|x\rangle.$ The latter remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if and only if $\zeta^{L}=1$.
Because of the U(1) symmetry of the Hamiltonian in Eq. (38), the plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (16) is actually an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$. The eigen energy is given by
$E(k)$ $=\frac{|e^{ik}-\zeta|^{2}}{1+|\zeta|^{2}}=\frac{\big{|}e^{i(k-k_{0})}-|\zeta|\big{|}^{2}}{1+|\zeta|^{2}}$
$=\frac{1-|\zeta|^{2}}{1+|\zeta|^{2}}+\frac{4|\zeta|}{1+|\zeta|^{2}}\sin^{2}((k-k_{0})/2),$ (44)
where $k_{0}$ is defined by $\zeta=|\zeta|e^{ik_{0}}$. Hence, the model is gapless and the conjecture holds when $|\zeta|=1$, regardless of $\zeta^{L}=1$ or not.
III.3 Degeneracy in gapped cases
Although our main focus is on gapless models, let us discuss the ground state degeneracy of gapped cases for the model in Eq. (14). This result will be used in our discussion for higher dimensional models in Sec. IV.
Let us start with the rank 1 model. In the $\gamma^{2}\neq\beta^{2}$ case, the formula Eq. (32) implies that $N_{\mathrm{deg}}=2$, because $m_{\psi}^{L}$ cannot be proportional to $\mathbbm{1}$ when the absolute values of the diagonal elements $\alpha-i\beta$ and $-\alpha+i\gamma$ do not agree. The $\alpha=\beta=\gamma=0$ case is exceptional. In this case $|\psi\rangle=\pm|11\rangle$ is a symmetric product state and any product state with no consecutive $1$’s is a ground state. Hence, the ground state degeneracy under PBC with $L\geq 3$ is given by the Lucas number Wikipedia article
$N_{\mathrm{deg}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2},$ (45)
which increases exponentially with the system size $L$. Here $\lfloor x\rfloor$ represents the greatest integer less than or equal to $x\in\mathbb{R}$. We append the derivation in Appendix C.
Next, let us discuss the rank 2 case. When $|\zeta|\neq 1$, the ground state of $\hat{H}^{\mathrm{pbc}}$ composed of the projector in Eq. (38) is unique ($N_{\mathrm{deg}}=1$) because $\zeta^{L}$ cannot be 1. According to Ref. Bravyi and Gosset (2015), $N_{\mathrm{deg}}\leq 2$ holds even for other rank 2 models under OBC. The min-max principle in Eq. (12) implies the same under PBC.
In summary, for translation-invariant $s=1/2$ chains with nearest neighbor interactions, the ground state degeneracy in the gapped cases is bounded as
$N_{\mathrm{deg}}\leq 2,$ (46)
except for the rank 1 model with $\alpha=\beta=\gamma=0$.
Figure 1: The illustration of the 2D model in Eq. (47) and its decomposition into 1D models.
IV 2D models with nearest-neighbor interactions
In this section, we generalize the results in the previous section to two and three dimensional models using the min-max principle.
We consider a square lattice of $s=1/2$ spins and assume the Hamiltonian of the form:
$\hat{H}^{\mathrm{2D}}\coloneqq\sum_{x,y=1}^{L}\big{(}\hat{Q}_{(x,y),(x+1,y)}+\hat{Q}_{(x,y),(x,y+1)}\big{)},$ (47)
where $\hat{Q}_{(x,y),(x^{\prime},y^{\prime})}$ is a projector acting on two spins at $(x,y)$ and $(x^{\prime},y^{\prime})$ [Fig. 1 (a)]. We impose PBC and assume the translation invariance in both $x$ and $y$ directions. Our goal is to show that when $\hat{H}^{\mathrm{2D}}$ is frustration free and gapless, there exists a variational state with properties in Eqs. (8) and (9).
Let us decompose the 2D Hamiltonian into 1D chains in two different ways.
$\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}.$ (48)
The first choice is the simple one, decomposing the system into decoupled chains as illustrated in Fig. 1 (b) and (c):
$\hat{H}^{(1)}\coloneqq\sum_{y=1}^{L}\hat{H}_{y}^{\mathrm{1D}},\quad\hat{H}_{y}^{\mathrm{1D}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{(x,y),(x+1,y)},$ (49)
$\hat{H}^{(2)}\coloneqq\sum_{x=1}^{L}\hat{H}_{x}^{\mathrm{1D}},\quad\hat{H}_{x}^{\mathrm{1D}}\coloneqq\sum_{y=1}^{L}\hat{Q}_{(x,y),(x,y+1)}.$ (50)
For each $y$, $\hat{H}_{y}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $x$ axis; similarly, for each $x$, $\hat{H}_{x}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $y$ axis. The second choice is illustrated in Fig. 1 (d) and (e), in which the chains are connected into one piece.
Even when these 1D Hamiltonians are frustration-free, whether the total Hamiltonian $\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}$ remains frustration-free or not is, in general, nontrivial. However, this issue can be easily gone around in our case. When either $\hat{H}^{\prime}{}^{(1)}$ or $\hat{H}^{\prime}{}^{(2)}$ is gapped with a finite ground state degeneracy, the min-max principle suggests that $\hat{H}^{\mathrm{2D}}$ is also gapped, because $\hat{H}^{\prime}{}^{(2)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(1)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(2)}$ are positive semi-definite  [^1] This argument cannot be directly applied to the simpler decomposition $\hat{H}^{(1)}+\hat{H}^{(2)}$, because if $N_{\mathrm{deg}}\geq 2$ for each $\hat{H}_{y}^{\mathrm{1D}}$, then the ground state degeneracy of $\hat{H}^{(1)}$ becomes $N_{\mathrm{deg}}^{L}$, which grows exponentially with the system size and the min-max principle becomes silent. For the same reason, the case of $\alpha=\beta=\gamma=0$ in the rank 1 model should be treated separately, but this case is trivially gapped in any spatial dimension. . Hence, we can restrict ourselves to the case where both $\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(2)}$ are gapless, implying that $\hat{H}^{(1)}$ and $\hat{H}^{(2)}$ are also gapless. Furthermore, as we have seen in Sec. III, when the Hamiltonian for an $s=1/2$ spin chain with nearest-neighbor interaction is gapless, the fully polarized state $|{\Phi_{0}}\rangle$ in Eq. (15) is a common zero-energy ground state among all models. Hence, the 2D version of the fully polarized state
$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x,y=1}^{L}|0\rangle_{(x,y)}$ (51)
is a simultaneous ground state of all terms in $\hat{H}^{\mathrm{2D}}$, implying that $\hat{H}^{\mathrm{2D}}$ is frustration-free.
To construct a variational state for low-energy excitations, let us define the plane-wave state of a single spin flip by
$|\Phi_{(k_{x},k_{y})}\rangle\coloneqq\frac{1}{L}\sum_{x,y=1}^{L}e^{i(k_{x}x+k_{y}y)}\hat{s}_{(x,y)}^{-}|{\Phi_{0}}\rangle.$ (52)
The variational energy is given by
$E(k_{x},k_{y})$ $\coloneqq\langle\Phi_{(k_{x},k_{y})}|\hat{H}^{\mathrm{2D}}|\Phi_{(k_{x},k_{y})}\rangle$
$=\sum_{y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{y}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
$\quad+\sum_{x=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{x}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
$=E(k_{x})+E(k_{y}),$ (53)
where we used the relation
$E(k_{x})=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik_{x}(x-x^{\prime})}\langle x^{\prime},y|\hat{H}_{y}^{\mathrm{1D}}|x,y\rangle,$ (54)
$E(k_{y})=\frac{1}{L}\sum_{y,y^{\prime}=1}^{L}e^{ik_{y}(y-y^{\prime})}\langle x,y^{\prime}|\hat{H}_{x}^{\mathrm{1D}}|x,y\rangle.$ (55)
Therefore,
$E(k_{x},k_{y})=O\big{(}(k_{x}-k_{0})^{2},(k_{y}-k_{0})^{2}\big{)}$ (56)
The discussion above can be readily extended to the cubic lattice in three dimensions. The variational energy of the plane-wave state $|\Phi_{(k_{x},k_{y},k_{z})}\rangle\coloneqq L^{-3/2}\sum_{x,y,z=1}^{L}e^{i(k_{x}x+k_{y}y+k_{z}z)}\hat{s}_{(x,y,z)}^{-}|{\Phi_{0}}\rangle$ is simply given by $E(k_{x})+E(k_{y})+E(k_{z})$.
Finally, let us discuss whether this analysis can be generalized to other form of lattices. For example, the nearest-neighbor interaction on the triangular lattice contains a diagonal interaction
$\sum_{x,y=1}^{L}\hat{Q}_{(x,y),(x+1,y-1)}.$ (57)
More generally, when an interaction $\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}$ among the spins at $(x,y)$ and $(x+d_{x},y+d_{y})$ is added, the variational energy obtains a term
$E(k_{x}d_{x}+k_{y}d_{y})$
$=\sum_{x,y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}|\Phi_{(k_{x},k_{y})}\rangle.$ (58)
Hence, the variational state generically becomes gapped unless $E(k_{0}(d_{x}+d_{y}))=0$. When $k_{0}=0$ (for example, in the ferromagnetic Heisenberg model), this condition can be easily satisfied, but otherwise some fine-tuning is required.
V 1D examples with next nearest-neighbor interactions
In Sec. III, we showed that frustration-free Hamiltonians for $s=1/2$ spin chains with nearest-neighbor interactions have a common ground state and a low-energy variational state. This observation was the key in demonstrating the softness of the dispersion relation in higher dimensional models in Sec. IV.
Table 1: Comparison of gapless frustration-free Hamiltonians for $s=1/2$ chain. $|{\Phi_{0}}\rangle$ [(15)] is the fully polarized state and $|{\Psi_{k}}\rangle$ [Eq. (16)] is the plane-wave state of single spin flip.
Interaction type Range $|{\Phi_{0}}\rangle$ is a GS $|{\Psi_{k}}\rangle$ is gapless Variational energy
Nearest neighbor models with rank 1: (19), (20) with $\gamma=\pm\beta$ 1 $O\big{(}(k-k_{0})^{2}\big{)}$
Nearest neighbor models with rank 2: (38) with $|\zeta|=1$ 1 $O\big{(}(k-k_{0})^{2}\big{)}$
Fredkin spin chain: (63) 2 $O\big{(}k^{2}\big{)}$
GHZ uncle Hamiltonian: (67) 2 $O\big{(}k^{2},L^{-2}\big{)}$
XYZ MG model: (78), (82) with $\theta\neq 0$ 2 $O\big{(}k^{2})$
In this section, we discuss $s=1/2$ spin chains with longer-range interactions. We assume that the Hamiltonian has the following form
$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1,x+2},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-2}\hat{Q}_{x,x+1,x+2}.$ (59)
Here, $\hat{Q}_{x,x+1,x+2}$ is a projector acting on the three spins at $x$, $x+1$, and $x+2$. The interaction range of the Hamiltonian is 2. Unlike the only nearest-neighbor interaction case in the previous section, our discussion here is based on examples and hence is not comprehensive. However, these examples clearly demonstrate difficulties in the direct extension of our results to more general cases. Namely, the fully polarized state $|{\Phi_{0}}\rangle$ may not be a ground state and the plane-wave state $|\Psi_{k}\rangle$ may not be a low-energy state. We instead construct a low-energy variational state which does not take the form $\sum_{x}e^{ikx}\hat{O}_{x}|\Phi_{0}\rangle$. We summarize our results in Table 1.
V.1 W state
Let us start with the uncle Hamiltonian constructed by the fully polarized state and the W state $|W\rangle\coloneqq(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ Fernández-González et al. (2015). The kernel of the Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}=\mathrm{span}\big{\{}|000\rangle,|W\rangle\big{\}}.$ (60)
The spin representation of the Hamiltonian is found to be
$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{3}{4}\hat{\mathbbm{1}}-\frac{\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}+\hat{s}_{x+2}^{z}}{3}+2\hat{s}_{x}^{z}\hat{s}_{x+1}^{z}\hat{s}_{x+2}^{z}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+1}^{z}\Big{)}\hat{\bm{s}}_{x+2}\cdot\hat{\bm{s}}_{x}$
$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+2}^{z}\Big{)}\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (61)
The fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of a single spin flip $|\Psi_{0}\rangle$ are the two ground states of this model regardless of the boundary condition. The state $|{\Psi_{k}}\rangle$ in Eq. (16) is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ as expected by the U(1) symmetry. The eigen energy is given by
$E(k)$ $=\big{|}\hat{Q}_{1,2,3}(|100\rangle+e^{ik}|010\rangle+e^{2ik}|001\rangle)\big{|}^{2}$
$=\big{|}\hat{Q}_{1,2,3}((e^{ik}-1)|010\rangle+(e^{2ik}-1)|001\rangle)\big{|}^{2}$
$=2(1-\cos k)\big{|}\hat{Q}_{1,2,3}(|010\rangle+(e^{ik}+1)|001\rangle)\big{|}^{2}$
$=\frac{8}{3}(2+\cos k)\sin^{2}(k/2).$ (62)
As a related model, let us discuss the bulk Hamiltonian for the Fredkin spin chain Salberger and Korepin (2017):
$\hat{Q}_{x,x+1,x+2}^{\prime}$ $\coloneqq\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}\Big{)}$
$\quad+\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}\Big{)}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+2}^{z}\Big{)}.$ (63)
The kernel of this Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}$ $=\mathrm{span}\Big{\{}|000\rangle,|100\rangle,|001\rangle+|010\rangle,$
[…]
Since $\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}\supset\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$, the min-max principle implies that the Fredkin spin chain is gapless with dispersion equal to or softer than the one implied by Eq. (62). Indeed, $|{\Psi_{k}}\rangle$ is an exact eigenstate with eigenenergy
$E^{\prime}(k)$ $=1-\cos k=2\sin^{2}(k/2).$ (65)
In these two examples, the fully polarized state $|{\Phi_{0}}\rangle$ is one of the ground states and the plane-wave state $|\Psi_{k}\rangle$ is a low-energy state with the dispersion $E(k)=O(k^{2})$.
V.2 Greenberger–Horne–Zeilinger uncle Hamiltonian
Next, we discuss the uncle Hamiltonian for the Greenberger–Horne–Zeilinger (GHZ) state Fernández-González et al. (2015). The kernel of the Hamiltonian is
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
$=\mathrm{span}\Big{\{}|000\rangle,|111\rangle,|100\rangle+|110\rangle,|001\rangle+|011\rangle\Big{\}}.$ (66)
The spin representation the Hamiltonian is
$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+1}^{z}(\hat{s}_{x}^{z}+\hat{s}_{x+2}^{z})$
$\quad-2\hat{s}_{x+1}^{x}\Big{(}\frac{1}{4}-\hat{s}_{x}^{z}\hat{s}_{x+2}^{z}\Big{)}.$ (67)
To write down ground states and low-energy excitations of this model, let us define
$|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle,\quad|\bar{n}\rangle\coloneqq|\underbrace{1\cdots 1}_{L-n}\underbrace{0\cdots 0}_{n}\rangle.$ (68)
for $n=1,2,\cdots,L-1$. For example, $|n\rangle$ satisfies
$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
$=-\frac{1}{2}(|n+1\rangle+|n-1\rangle+\hat{T}|n+1\rangle+\hat{T}^{-1}|n-1\rangle)$ (69)
for $n=2,3,\cdots,L-2$, and
$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
=$=\begin{cases}-\frac{1}{2}(|2\rangle+\hat{T}|2\rangle)&(n=1),\\ -\frac{1}{2}(|L-2\rangle+\hat{T}^{-1}|L-2\rangle)&(n=L-1)\end{cases}$ (70)
for $n=1$ and $L-1$. Based on these expressions, the ground states of $\hat{H}^{\mathrm{obc}}$ are found to be
$|0\cdots 0\rangle\,(=|\Phi_{0}\rangle),|1\cdots 1\rangle,\sum_{n=1}^{L-1}|n\rangle,\sum_{n=1}^{L-1}|\bar{n}\rangle.$ (71)
The latter two are not consistent with the PBC and are hence not ground states of $\hat{H}^{\mathrm{pbc}}$.
Although the fully-polarized state $|{\Phi_{0}}\rangle$ is one of the ground states, the simple variational state $|{\Psi_{k}}\rangle$ is not a low-energy state:
$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=2.$ (72)
We instead consider a plane-wave state of two domain walls:
$|\Psi_{k,\ell}\rangle\coloneqq\frac{\sqrt{2}}{L}\sum_{n=1}^{L-1}\sum_{m=0}^{L-1}e^{ikm}e^{-ikn/2}\sin(\tfrac{\pi n\ell}{L})\hat{T}^{m}|n\rangle$ (73)
for $\ell=1,2,\cdots,L-1$. Using $\hat{T}|\Psi_{k,\ell}\rangle=e^{-ik}|\Psi_{k,\ell}\rangle$ and $[\hat{T},\hat{H}^{\mathrm{pbc}}]=0$, one can readily show that $|\Psi_{k,\ell}\rangle$ is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ with the eigenvalue
$E_{\ell}(k)$ $=2[1-\cos(k/2)\cos(\tfrac{\pi\ell}{L})]$
$=4\sin^{2}(\tfrac{\pi\ell}{2L})+4\cos(\tfrac{\pi\ell}{L})\sin^{2}(k/4).$ (74)
The first term in the last expression is $O(L^{-2})$, which vanishes in the thermodynamic limit. The second term is $O(k^{2})$.
Despite the presence of gapless excitations, in this model, the correlation function of the form
$\langle\Phi_{0}|\hat{\mathcal{O}}_{\bm{x}}^{\dagger}(\hat{\mathbbm{1}}-\hat{G})\hat{\mathcal{O}}^{\prime}_{\bm{y}}|\Phi_{0}\rangle$ (75)
decays exponentially with the system size for any choice of local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}^{\prime}_{\bm{y}}$. Here $\hat{G}$ is the projector onto the ground states of $\hat{H}^{\mathrm{pbc}}$. This is because both of the two ground states of $\hat{H}^{\mathrm{pbc}}$ (the first two states in Eq. (71)) are product states.
V.3 XYZ MG model
As our last example, let us discuss variations of the MG model. The local Hamiltonian of the original model is given by
$\hat{Q}_{x,x+1,x+2}$
$\coloneqq\frac{1}{2}\hat{\mathbbm{1}}+\frac{2}{3}(\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}+\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}+\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+2}).$ (76)
Any singlet state $|s\rangle_{x,x+1}\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ belongs to the kernel of this Hamiltonian.
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
$=\mathrm{span}\Big{\{}|010\rangle-|001\rangle,|100\rangle-|010\rangle,$
[…]
The extension of this model to the XYZ coupling
$\hat{Q}_{x,x+1,x+2}^{\prime}$
$\coloneqq J_{0}\hat{\mathbbm{1}}+\sum_{a=x,y,z}J_{a}(\hat{s}_{x}^{a}\hat{s}_{x+1}^{a}+\hat{s}_{x+1}^{a}\hat{s}_{x+2}^{a}+\hat{s}_{x}^{a}\hat{s}_{x+2}^{a})$ (78)
is still frustration free for some ranges of $J_{x}$, $J_{y}$, $J_{z}$ Sriram Shastry and Sutherland (1981); Shastry and Sutherland (1981); Gerhardt et al. (1998); Xu et al. (2021); Palle and Benton (2021). In particular, when
$J_{x}J_{y}+J_{y}J_{z}+J_{z}J_{x}=0,$ (79)
the model exhibits additional ground state degeneracy. Following Ref. Palle and Benton (2021), we set $J_{0}=\frac{1}{4}$ and
$J_{x}$ $=\frac{1}{3}-\frac{2}{3}\cos\Big{(}2\theta-\frac{2\pi}{3}\Big{)},$ (80)
$J_{y}$ $=\frac{1}{3}-\frac{2}{3}\cos\Big{(}2\theta+\frac{2\pi}{3}\Big{)},$ (81)
$J_{z}$ $=\frac{1}{3}-\frac{2}{3}\cos 2\theta$ (82)
with $-\pi/2\leq\theta\leq\pi/2$ so that the condition in Eq. (79) is satisfied and $\hat{Q}_{x,x+1,x+2}^{\prime}$ becomes a projector. The kernel of $\hat{Q}_{x,x+1,x+2}^{\prime}$ is given by
$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}=\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}\oplus\mathrm{span}\big{\{}|\phi\rangle,|\bar{\phi}\rangle\big{\}},$ (83)
where  [^2] The presence of $|\bar{\phi}\rangle$ in the kernel is not relevant. Hence, the rank of $\hat{Q}_{x,x+1,x+2}$ can be three rather than two.
$|\phi\rangle=\cos\theta\,|000\rangle+\frac{\sin\theta}{\sqrt{3}}(|011\rangle+|101\rangle+|110\rangle),$ (84)
$|\bar{\phi}\rangle=\cos\theta\,|111\rangle+\frac{\sin\theta}{\sqrt{3}}(|100\rangle+|010\rangle+|001\rangle).$ (85)
Unless $\theta=0$, the fully polarized state $|{\Phi_{0}}\rangle$ is not a ground state of the system. The $\theta=\pm\pi/4$ cases are unitary equivalent to the model discussed in Ref. Saito and Hotta (2024). The plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (16) is not a low-energy state unless $\theta=0$; indeed, we find
$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\frac{\cos^{2}\theta}{3}(1+2\cos k)^{2}+(L-3)\sin^{2}\theta.$ (86)
To construct a low-energy excitation, we assume $L$ is an odd integer greater than three. We introduce a plane-wave state of the domain wall state
$|\tilde{\Psi}_{k}\rangle=\frac{1}{\sqrt{Ln_{k}}}\sum_{m=0}^{L-1}e^{ikm}\hat{T}^{m}|D\rangle,$ (87)
$|D\rangle\coloneqq|\phi\rangle_{1,2,3}|s\rangle_{4,5}|s\rangle_{6,7}\cdots|s\rangle_{L-1,L},$ (88)
where $n_{k}\coloneqq 1-(-2)^{-\frac{L-5}{2}}\cos k$ is the normalization factor that converges to $1$ in the large $L$ limit. Using the translation symmetry of $\hat{H}$, we find
$\langle\tilde{\Psi}_{k}|\hat{H}|\tilde{\Psi}_{k}\rangle$ $=\frac{\big{\|}\hat{Q}_{2,3,4}(|D\rangle+\hat{T}^{2}|D\rangle)\big{\|}^{2}}{n_{k}}=\frac{2\sin^{2}k}{3n_{k}}.$ (89)
This is $O(k^{2})$, which is consistent with the result of Ref. Palle and Benton (2021) on Kagome lattice. Hence, our conjecture still holds in this example as well, despite the fact that $|{\Phi_{0}}\rangle$ is not a ground state and $|{\Psi_{k}}\rangle$ is not gapless.
VI Conclusion
In this paper, we proposed two conjectures regarding frustration-free systems. The first conjecture posited that no finite-size gap opens between the degenerate ground states of frustration-free systems. While this has been implicitly assumed in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016), we pointed out the necessity of being cautious as it is actually a non-trivial assumption.
The second conjecture stated that in a gapless frustration-free system with translational symmetry, the dispersion relation of low-energy excitations near a certain wave number $\bm{k}_{0}$ can be bounded above by a quadratic dispersion, i.e., $O\big{(}|\bm{k}-\bm{k}_{0}|^{2},L^{-2}\big{)}$. We proved this conjecture for the case of an $s=1/2$ spin chain with nearest-neighbor interactions by combining the result of a seminal work by Bravyi and Gosset Bravyi and Gosset (2015) and the min-max theorem. Furthermore, we extended these results to two- and three-dimensional models on the cubic lattice. However, it became clear that this method could not be directly extended to cases where three or more neighboring spins interact simultaneously. Extending our results to cases where the Hilbert space dimension at each site is larger than two or to models with longer-range interactions constitutes important future work.
Finally, there is a related, more general conjecture that the finite-size gap $\epsilon=E_{\tilde{N}_{\mathrm{deg}}+1}$ for a gapless frustration-free Hamiltonian can be bounded above by $O(L^{-2})$ regardless of the presence or absence of translational symmetry or the details of the boundary conditions. In fact, in the accompanying paper Masaoka et al. (2024), we prove this statement for critical frustration-free systems in which a ground-state correlation function shows a power-law behavior. However, such an argument is not applicable, for example, to the uncle Hamiltonian for the GHZ state since all correlation functions decays exponentially as discussed in Sec. V.2.
Acknowledgements.
We thank Hosho Katsura for useful discussions. The work of H.W. is supported by JSPS KAKENHI Grant No. JP24K00541. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S.
Appendix A Proof of the inequality (11)
Let $\mathcal{M}_{j}$ be an arbitrary $j$ dimensional subspace of the entire Hilbert space $\mathcal{H}_{D}$. We maximize the energy expectation value $\langle\Phi|\hat{H}|\Phi\rangle$ by varying the normalized state $|\Phi\rangle$ belonging to $\mathcal{M}_{j}$. We then minimize the maximum value $\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}|\Phi\rangle$ by varying the subspace $\mathcal{M}_{j}$ of $\mathcal{H}_{D}$. By definition, the minimum value $E_{j}$ is achieved when $\mathcal{M}_{j}$ is spanned by the $j$ eigenvectors corresponding to the eigenvalues $E_{1},E_{2},\cdots,E_{j}$. Hence, we obtain an expression
$E_{j}=\min_{\mathcal{M}_{j}\subset\mathcal{H}_{D}}\Big{(}\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}|\Phi\rangle\Big{)}.$ (90)
The same argument leads to
$E_{j}^{\prime}=\min_{\mathcal{M}_{j}\subset\mathcal{H}_{D}}\Big{(}\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle\Big{)}.$ (91)
Finally, by the definition of $\hat{V}=\hat{H}^{\prime}-\hat{H}\geq 0$, we have
$\langle\Phi|\hat{H}^{\prime}|\Phi\rangle\geq\langle\Phi|\hat{H}|\Phi\rangle$ (92)
for any state $|\Phi\rangle$ in $\mathcal{H}_{D}$. In particular, if $\langle\Phi|\hat{H}|\Phi\rangle$ and $\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ are maximized by $|\Phi\rangle_{*}\in\mathcal{M}_{j}$ and $|\Phi\rangle_{*}^{\prime}\in\mathcal{M}_{j}$, respectively, we have
$\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ $=\langle\Phi_{*}^{\prime}|\hat{H}^{\prime}|\Phi_{*}^{\prime}\rangle$
$\geq\langle\Phi_{*}|\hat{H}^{\prime}|\Phi_{*}\rangle$
$\geq\langle\Phi_{*}|\hat{H}|\Phi_{*}\rangle=\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}|\Phi\rangle.$ (93)
This relation holds for any $\mathcal{M}_{j}\subset\mathcal{H}_{D}$. Hence, we arrive at the inequality (11).
Appendix B Derivation of Eq. (20)
Let us consider a local unitary that maps $|\psi\rangle$ to $|\psi^{\prime}\rangle=\hat{U}\otimes\hat{U}|\psi\rangle$. In general, $\hat{U}$ can be parametrized by $u_{1},u_{2}\in\mathbb{C}$ ($|u_{1}|^{2}+|u_{2}|^{2}=1$) and $\theta\in[0,2\pi)$ as
$\hat{U}|0\rangle=e^{i\theta}(u_{1}|0\rangle+u_{2}|1\rangle),$ (94)
$\hat{U}|1\rangle=e^{i\theta}(-u_{2}^{*}|0\rangle+u_{1}^{*}|1\rangle),$ (95)
The inverse of $\hat{U}$ reads
$\hat{U}^{\dagger}|0\rangle=e^{-i\theta}(u_{1}^{*}|0\rangle-u_{2}|1\rangle),$ (96)
$\hat{U}^{\dagger}|1\rangle=e^{-i\theta}(u_{2}^{*}|0\rangle+u_{1}|1\rangle).$ (97)
Hence
$m_{\psi^{\prime}}$ $=\begin{pmatrix}\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|01\rangle&\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|11\rangle\\ -\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|00\rangle&-\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|10\rangle\end{pmatrix}$
$=e^{-2i\theta}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)\end{pmatrix}$
$=e^{-2i\theta}\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|1\rangle&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|1\rangle\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|0\rangle&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|0\rangle\end{pmatrix}$
$=e^{-2i\theta}\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}\begin{pmatrix}u_{1}^{*}&u_{2}^{*}\\ -u_{2}&u_{1}\end{pmatrix}$
$=e^{-2i\theta}Um_{\psi}U^{\dagger}.$ (98)
Therefore, by choosing $U=\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}$ and $\theta$ properly, one gets the form in Eq. (25), which is equivalent to the parametrization in Eq. (20).
Appendix C The case of $\det m_{\psi}=0$.
Here we discuss the case $\alpha=\beta=0$ in the rank 1 model. (The $\alpha=\gamma=0$ case can be treated in the same way.) In this case, $|\psi\rangle$ is not entangled because it can be written as a product state
$|\psi\rangle=|1\rangle\otimes|v^{\perp}\rangle,$ (99)
where $|v^{\perp}\rangle=i\gamma|0\rangle+\delta|1\rangle$ with $\gamma^{2}+\delta^{2}=1$. According to Ref. Bravyi and Gosset (2015), excitations under OBC are gapped in this case, which implies that excitations under PBC are also gapped.
When $\gamma\neq 0$, orthogonal ground states under OBC are product states with a single domain-wall:
$|\underbrace{0\cdots 0}_{n-1}v^{\perp}\underbrace{v\cdots v}_{L-n}\rangle\quad(n=1,\cdots L),$ (100)
where $|v\rangle=\delta|0\rangle+i\gamma|1\rangle$ is orthogonal to $|v^{\perp}\rangle$. Translation-invariant product states $|v\cdots v\rangle$ and $|0\cdots 0\rangle$ are also ground states but only the former is orthogonal to the domain-wall states and the latter is not linearly independent. In contrast, under PBC, domain wall states violate the boundary term and the two translation-invariant product states give the ground states. Thus the formula in Eq. (32) is applicable to this case.
On the other hand, the $\alpha=\beta=\gamma=0$ case is exceptional. In this case $|\psi\rangle=\pm|11\rangle$ is a symmetric product state and any product state with no consecutive $1$’s is a ground state of $\hat{H}^{\mathrm{obc}}$ and $\hat{H}^{\mathrm{pbc}}$. Hence,
$N_{\mathrm{deg}}^{\mathrm{obc}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n},$ (101)
$N_{\mathrm{deg}}^{\mathrm{pbc}}=N_{\mathrm{deg}}^{\mathrm{obc}}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2},$ (102)
where $\lfloor x\rfloor$ represents the greatest integer less than or equal to $x\in\mathbb{R}$. The second term in $N_{\mathrm{deg}}^{\mathrm{pbc}}$ represents the number of valid product states under OBC which start with $|1\rangle_{x=1}$ and end with $|1\rangle_{x=L}$. The excitation gap is exactly 1.

### v2

@@
 Recent case-by-case studies revealed that the dispersion of low energy excitations in gapless frustration-free Hamiltonians is often quadratic or softer.
-In this work, we argue that this is actually a general property of such systems.
-By combining a previous study by Bravyi and Gosset and the min-max principle, we prove this hypothesis for models with local Hilbert spaces of dimension two that contains only nearest-neighbor interactions on cubic lattice.
+In this worl, we argue that this is actually a general property of such systems.
+By combining the previous study by Bravyi and Gosset and the min-max principle, we prove this hypothesis for models with local Hilbert spaces of dimension two that contains only nearest-neighbor interactions on cubic lattice.
 This may be understood as a no-go theorem realizing gapless phases with linearly dispersive excitations in frustration-free Hamiltonians.
@@
 It is hence common practice to analyze the system of interest using the simplest possible model that belongs to the same phase.
-Frustration-free systems, in particular, are relatively easy to handle, and their exact ground states and excited states are sometimes accessible.
-For these reasons, frustration-free systems have been widely used as representative models of various phases of matter.
-Therefore, elucidating the general properties of frustration-free systems and understanding the limits of the phases that can be realized by frustration-free Hamiltonians hold significant importance.
-Numerous frustration-free systems posses an excitation gap.
-Notable examples include the Majumdar–Ghosh (MG) model, known as a toy model realizing spontaneous breaking of discrete translational symmetry, the Affleck–Kennedy–Lieb–Tasaki (AKLT) model, representing a symmetry-protected topological phase, and the Kitaev Toric code model, serving as the canonical model for topologically ordered phases Auerbach (2012); Tasaki (2020).
+Frustration-free (FF) systems, in particular, are relatively easy to handle, and their exact ground states and excited states are sometimes accessible.
+For these reasons, FF systems have been widely used as representative models of various phases of matter.
+Therefore, elucidating the general properties of FF systems and understanding the limits of the phases that can be realized by FF Hamiltonians hold significant importance.
+Numerous FF systems possess an excitation gap.
+Notable examples include the Majumdar–Ghosh (MG) model, known as a toy model realizing spontaneous breaking of discrete translational symmetry, the Affleck–Kennedy–Lieb–Tasaki model, representing a symmetry-protected topological phase, and the Kitaev toric code model, serving as the canonical model for topologically ordered phases Auerbach (2012); Tasaki (2020).
 The parent Hamiltonians of Matrix Product States (MPS) also feature an excitation gap Fannes et al.
 (1992).
-There are also frustration-free Hamiltonians with gapless excitations.
+There are also FF Hamiltonians with gapless excitations.
 For example, the uncle Hamiltonian of MPS Fernández-González et al.
@@
 (2023) are gapless.
-Through recent investigation of these models, the general properties of frustration-free systems have gradually become clearer.
-In this paper, we discuss two conjectures regarding frustration-free systems.
-The first is about the finite-size gap between degenerate ground states.
-Generally, even if the ground states are degenerate in the thermodynamic limit, their energy eigenvalues in a finite-size system do not precisely match due to the finite-size gap.
-However, in frustration-free systems, the ground states are exactly degenerate even before taking the thermodynamic limit.
-In other words, we can judge whether a frusration-free system is gapped or not by looking at the limiting behavior of the difference between the ground state energy and the second lowest energy level.
-The second conjecture concerns the dispersion relation of low-energy excitations in frustration-free systems.
-Although gapless modes typically have a linear dispersion, it is known that the ferromagnetic Heisenberg model, which is a representative frustration-free Hamiltonian, have a spin wave excitation with a quadratic dispersion Auerbach (2012).
-In fact, our second conjecture asserts that gapless frustration-free systems generally have a quadratic or softer dispersion relation.
-Although there are several arguments supporting this statement in literature (for example, see Refs.
-Palle and Benton, 2021; Ogunnaike et al., 2023; Watanabe et al., 2023; Ren et al., 2024), they are applicable only to the case where the low-energy excitations are obtained by applying a sum of local operators to a ground state.
-We provide a general proof for spin-1/2 models with nearest-neighbor interaction on cubic lattice based on the result of Bravyi and Gosset Bravyi and Gosset (2015).
+Through their recent investigations, the general properties of FF systems have gradually become clearer.
+In this worl, we discuss two conjectures regarding FF systems.
+The first one is about the finite-size splitting between degenerate ground states.
+Generally, even if the ground states are degenerate in the thermodynamic limit, their energy eigenvalues in a finite system do not match due to the finite-size splitting.
+However, in FF systems, the ground states are exactly degenerate even before taking the thermodynamic limit.
+In other words, whether a frusration-free system is gapped or not can be judged by looking at the limiting behavior of the difference between the ground state energy and the second lowest energy level.
+The second conjecture concerns the dispersion relation of low-energy excitations in FF systems.
+Although gapless modes typically have a linear dispersion, it is known that the ferromagnetic Heisenberg model, a representative FF Hamiltonian, have a spin wave excitation with a quadratic dispersion $E(k)=2\sin^{2}(k/2)$ Auerbach (2012).
+In fact, our second conjecture asserts that gapless FF systems always have a quadratic or softer dispersion relation.
+Although there are several arguments supporting this statement in literature Palle and Benton (2021); Ogunnaike et al.
+(2023); Watanabe et al.
+(2023); Ren et al.
+(2024), they are applicable only to the case where the low-energy excitations are obtained by applying a sum of local operators to a ground state.
+We provide a general proof for spin-1/2 models with nearest-neighbor interaction on cubic lattice based on Ref.
+Bravyi and Gosset, 2015.
 We also discuss examples with longer-range interactions and show that gapless excitations with quadratic dispersion may not be obtained by a sum of local operators.
 II Overview
-II.0.1 Setting and definitions
-In this work, we consider Hamiltonians defined on a finite $d$-dimensional lattice $\Lambda$:
-$\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{H}_{\bm{r}}.$ (1)
-The local Hilbert space on each lattice site $\bm{r}\in\Lambda$ is assumed to be finite dimensional, which is denoted by $D_{0}$.
+II.1 Setting and definitions
+In this work, we consider Hamiltonians
+$\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{H}_{\bm{r}}$ (1)
+defined on a finite $d$-dimensional lattice $\Lambda$.
+The local Hilbert space on each lattice site $\bm{r}\in\Lambda$ is finite dimensional, which is denoted by $D_{0}$.
 Let $|\Lambda|$ be the number of lattice sites in $\Lambda$.
@@
 We assume that the Hamiltonian is local in the sense that $\hat{H}_{\bm{r}}$ acts nontrivially on the sites within a finite distance $R$ from $\bm{r}$.
+The Hamiltonian is translation invariant if $\hat{T}_{\bm{a}}\hat{H}_{\bm{r}}=\hat{H}_{\bm{r}+\bm{a}}\hat{T}_{\bm{a}}$ for each lattice vector $\bm{a}$ and $\bm{r}\in\Lambda$.
+This is possible only when the periodic boundary condition (PBC) is imposed.
 Let us write eigenvalues of $\hat{H}$ as
@@
 The thermodynamic limit $|\Lambda|\to\infty$ is taken by a sequence of increasing system size.
-The system is said to be gapped if there exists an integer $N_{\mathrm{deg}}$ ($1\leq N_{\mathrm{deg}}<D$), which may depend on the system size, such that
-$\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}}-E_{1}\big{)}=0,$ (3)
-$\Delta\coloneqq\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}+1}-E_{N_{\mathrm{deg}}}\big{)}\neq 0.$ (4)
-Otherwise the system is gapless.
-The limit in Eq.
-(3) defines the excitation gap $\Delta$.
-When $N_{\mathrm{deg}}\geq 2$, the quantity $E_{N_{\mathrm{deg}}}-E_{1}$ is referred to as the finite-size gap among degenerate ground states.
+The system is said to be gapped if there exists an integer $N_{\mathrm{deg}}$ ($1\leq N_{\mathrm{deg}}<D$), which may depend on the system size, such that $\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}}-E_{1}\big{)}=0$ and $\lim_{|\Lambda|\to\infty}\big{(}E_{N_{\mathrm{deg}}+1}-E_{N_{\mathrm{deg}}}\big{)}\neq 0$.
+Otherwise the system is gapless, and the spectrum above the ground state is continuous.
+When $N_{\mathrm{deg}}\geq 2$, the quantity $E_{N_{\mathrm{deg}}}-E_{1}$ is referred to as the finite-size splitting among degenerate ground states.
 When the degeneracy originates from spontaneous breaking of a discrete symmetry or formation of a topological order, it generically decays exponentially with the system size.
 Even when $N_{\mathrm{deg}}$ converges to a finite number in the large $|\Lambda|$ limit, its value for a finite system may be ambiguous due to the possible system-size dependence.
-The Hamiltonian is translation invariant if there exists an operator $\hat{T}_{\bm{a}}$ for each lattice vector $\bm{a}$ such that $\hat{T}_{\bm{a}}\hat{H}_{\bm{r}}=\hat{H}_{\bm{r}+\bm{a}}\hat{T}_{\bm{a}}$ for all $\bm{r}\in\Lambda$.
-This is possible only when the periodic boundary condition (PBC) is imposed.
-The Hamiltonian is frustration-free if a ground state $|{\Phi_{0}}\rangle$ of $\hat{H}$ is a simultaneous ground state of every local Hamiltonian $\hat{H}_{\bm{r}}$ Tasaki (2020).
-In this case, without loss of generality and without modifying the ground states and low-energy excitations of the system, one may assume that each $\hat{H}_{\bm{r}}$ is a projector, i.e., $\hat{Q}_{\bm{r}}^{2}=\hat{Q}_{\bm{r}}$ and $\hat{Q}_{\bm{r}}|{\Phi_{0}}\rangle=0$ for all $\bm{r}\in\Lambda$.
+The Hamiltonian is FF if a ground state $|{\Phi_{0}}\rangle$ of $\hat{H}$ is a simultaneous ground state of every local Hamiltonian $\hat{H}_{\bm{r}}$ Tasaki (2020).
+In this case, without loss of generality nor modifying the ground states and low-energy excitations of the system, one may assume that each $\hat{H}_{\bm{r}}$ is a projector, i.e., $\hat{Q}_{\bm{r}}^{2}=\hat{Q}_{\bm{r}}$ and $\hat{Q}_{\bm{r}}|{\Phi_{0}}\rangle=0$ for all $\bm{r}\in\Lambda$.
 Under this choice, the ground state energy becomes $E_{1}=0$.
-II.0.2 Conjectures on frustration-free Hamiltonian
+II.2 Conjectures
 Let us state our conjectures one by one.
-The first conjecture is that, when the Hamiltonian is frustration-free and gapped, the finite-size gap among ground states precisely vanishes before taking the thermodynamic limit:
-$E_{N_{\mathrm{deg}}}=E_{1}.$ (5)
+II.2.1 Conjecture 1
+The first conjecture is that, when the Hamiltonian is FF and gapped, the finite-size splitting among ground states precisely vanishes before taking the thermodynamic limit; i.e., $E_{N_{\mathrm{deg}}}=E_{1}$.
 For example, in the AKLT model for spin $s=1$ chain under OBC, the four-fold ground state degeneracy due to the edge modes is exact.
 The four-fold topological degeneracy in the Kitaev toric code under PBC is also exact.
-This implies the following criterion for the excitation gap in frustration-free systems.
-Let $\tilde{N}_{\mathrm{deg}}$ ($1\leq\tilde{N}_{\mathrm{deg}}<D$) be the number of exact zero energy states, i.e.,
-$E_{\tilde{N}_{\mathrm{deg}}}=E_{1},\quad E_{\tilde{N}_{\mathrm{deg}}+1}\neq E_{1},$ (6)
-which may also depend on the system size (see, for example, Eq. (45) below). The Hamiltonian is gapped if and only if
-$\lim_{|\Lambda|\to\infty}\big{(}E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\big{)}>0.$ (7)
-Namely, one can set $N_{\mathrm{deg}}=\tilde{N}_{\mathrm{deg}}$.
-Actually this criterion has been used as the definition of the excitation gap in previous studies (for example, see Refs.
-Bravyi and Gosset, 2015; Gosset and Mozgunov, 2016; Gosset and Huang, 2016).
+This observation implies the following criterion for the excitation gap in FF systems.
+Let $\tilde{N}_{\mathrm{deg}}$ ($1\leq\tilde{N}_{\mathrm{deg}}<D$) be the number of exact zero energy states, i.e., $E_{\tilde{N}_{\mathrm{deg}}}=0$ and $E_{\tilde{N}_{\mathrm{deg}}+1}\neq 0$, which may also depend on the system size.
+The Hamiltonian is gapless if and only if
+$\lim_{|\Lambda|\to\infty}E_{\tilde{N}_{\mathrm{deg}}+1}=0.$ (3)
+Actually this criterion has been used as the definition of the excitation gap in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016).
 Note that $\tilde{N}_{\mathrm{deg}}$ is well-defined even for gapless systems unlike $N_{\mathrm{deg}}$.
-The second conjecture is that, when the Hamiltonian is translation-invariant, frustration-free, and gapless, there exists a family of variational states $|\Psi_{\bm{k}}\rangle$ that are orthogonal to all the ground states and satisfy
-$\hat{T}_{\bm{a}}|\Psi_{\bm{k}}\rangle=e^{-i\bm{k}\cdot\bm{a}}|\Psi_{\bm{k}}\rangle,$ (8)
+II.2.2 Conjecture 2
+The second conjecture is that, when the Hamiltonian is translation-invariant, FF, and gapless, there exists a family of variational states $|\Psi_{\bm{k}}\rangle$ that are orthogonal to all the ground states, are eigenstates of translation operator $\hat{T}_{\bm{a}}|\Psi_{\bm{k}}\rangle=e^{-i\bm{k}\cdot\bm{a}}|\Psi_{\bm{k}}\rangle$, and have a quadratic dispersion about the gapless point at $\bm{k}=\bm{k}_{0}$:
 $\langle\Psi_{\bm{k}}|\hat{H}|\Psi_{\bm{k}}\rangle=O\big{(}|\bm{k}-\bm{k}_{0}|^{2},L^{-2}\big{)}$ (4)
-for some $\bm{k}_{0}$. Namely, the low energy excitation of the system is not linearly dispersive about the gapless point at $\bm{k}=\bm{k}_{0}$.
 For systems under OBC, the result of Refs.
-Knabe, 1988; Anshu, 2020; Gosset and Mozgunov, 2016 gives us the bound $E_{\tilde{N}_{\mathrm{deg}}+1}-E_{1}\leq O(L^{-2})$.
+Knabe, 1988; Anshu, 2020; Gosset and Mozgunov, 2016; Lemm and Xiang, 2022 gives us the bound $E_{\tilde{N}_{\mathrm{deg}}+1}=O(L^{-2})$.
 Although this is consistent with our conjecture, it is not directly applicable to the PBC case Watanabe et al.
 (2023).
-II.0.3 Min-max principle
-Next, let us review the min-max principle, which supports our conjectures.
-We compare two Hamiltonians $\hat{H}$ and $\hat{H}^{\prime}$ acting on a $D$-dimensional Hilbert space.
+In Sec. III, we give a proof of this conjecture for $s=1/2$ spin chain with nearest neighbor interactions. We further generalize the results to higher dimensions in Sec. IV.
+II.3 Min-max principle
+Our discussions below are based on a mathematical theorem called the min-max principle (see, theorem A.7 of Ref.
+Tasaki, 2020).
+We compare two Hamiltonians $\hat{H}$ and $\hat{H}^{\prime}$ acting on the same $D$-dimensional Hilbert space.
 The eigenvalues of $\hat{H}$ and $\hat{H}^{\prime}$ are, respectively, denoted by $E_{j}$ and $E_{j}^{\prime}$ ($j=1,2,\cdots,D$) in the increasing order.
-Suppose that $\hat{V}\coloneqq\hat{H}^{\prime}-\hat{H}$ is positive semi-definite:
-$\hat{V}\geq 0.$ (10)
-That is, $\langle\Phi|\hat{V}|\Phi\rangle\geq 0$ for any state $|\Phi\rangle$. In other words, all the eigenvalues of $\hat{V}$ are nonnegative. In this setting, the min-max principle states that
-$E_{j}^{\prime}\geq E_{j}$ (11)
-for every $j=1,2,\cdots,D$ (see Theorem A.7 of Ref.
-Tasaki (2020)).
-This statement is trivial when the two Hamiltonians can be diagonalized by a common unitary operator, but it holds more generally.
-Since the proof is elementary, we review it in Appendix A.
-Two corollaries follow immediately from the min-max principle.
-Let us consider two frustration free Hamiltonians $\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}$ and $\hat{H}^{\prime}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}^{\prime}$.
+Suppose that $\hat{V}\coloneqq\hat{H}^{\prime}-\hat{H}$ is positive semi-definite $\hat{V}\geq 0$.
+That is, $\langle\Phi|\hat{V}|\Phi\rangle\geq 0$ for any state $|\Phi\rangle$.
+In this setting, the min-max principle states that
+$E_{j}^{\prime}\geq E_{j}\quad({}^{\forall}j=1,2,\cdots,D)$ (5)
+For readers’ convenience, we review the proof in Appendix A. This is trivial when the two Hamiltonians can be diagonalized by a common unitary operator, but it holds more generally.
+Two corollaries follow immediately from the theorem.
+Let us consider two FF Hamiltonians $\hat{H}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}$ and $\hat{H}^{\prime}=\sum_{\bm{r}\in\Lambda}\hat{Q}_{\bm{r}}^{\prime}$.
 Suppose that $\hat{V}=\hat{H}^{\prime}-\hat{H}$ is positive-semidefinite.
-Then the inequality in (11) implies that the number of zero-energy states of $\hat{H}$ cannot be smaller than that of $\hat{H}^{\prime}$:
-$\tilde{N}_{\mathrm{deg}}\geq\tilde{N}_{\mathrm{deg}}^{\prime}.$ (12)
-Furthermore,
-$\lim_{|\Lambda|\to\infty}E_{j}^{\prime}=0\quad\Rightarrow\quad\lim_{|\Lambda|\to\infty}E_{j}=0.$ (13)
-This relation suggests that if (i) $\hat{H}$ is gapped and (ii) $\tilde{N}_{\mathrm{deg}}$ is bounded by a system-size independent constant, then $\hat{H}^{\prime}$ is also gapped.
-This relation was previously used for interacting Kitaev chains in Refs.
-Katsura et al., 2015; Wouters et al., 2018.
-III 1D Examples with nearest-neighbor interactions
+Then the minimax principle implies that the number of zero-energy states of $\hat{H}$ cannot be smaller than that of $\hat{H}^{\prime}$: $\tilde{N}_{\mathrm{deg}}\geq\tilde{N}_{\mathrm{deg}}^{\prime}$.
+Furthermore, if (i) $\hat{H}$ is gapped and (ii) $\tilde{N}_{\mathrm{deg}}$ is bounded by a system-size independent constant, then $\hat{H}^{\prime}$ is also gapped.
+This is because, if $\hat{H}^{\prime}$ were gapless, an increasing number of eigenstates of $\hat{H}^{\prime}$ have lower energy than $E_{\tilde{N}_{\mathrm{deg}}+1}>0$ and eventually $E_{\tilde{N}_{\mathrm{deg}}+1}>E_{\tilde{N}_{\mathrm{deg}}+1}^{\prime}$ holds for a sufficiently large $L$.
+This relation was previously used for interacting Kitaev chains Katsura et al.
+(2015); Wouters et al.
+(2018).
+III 1D system with nearest-neighbor interactions
 […]
-$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-1}\hat{Q}_{x,x+1}.$ (14)
-Here, $\hat{Q}_{x,x+1}=\hat{Q}_{x,x+1}^{2}$ is a projector nontrivially acting on the spins at $x$ and $x+1$.
-The interaction range of this Hamiltonian is 1.
-The number of $+1$ eigenvalues of a projector is called the rank.
-The Hamiltonian $\hat{H}^{\mathrm{pbc}}$ has the translation symmetry $\hat{T}$ which satisfies $\hat{T}\hat{Q}_{x-1,x}=\hat{Q}_{x,x+1}\hat{T}$.
-The seminal work by Bravyi and Gosset Bravyi and Gosset (2015) showed that $\hat{H}^{\mathrm{obc}}$ can be gapless only when the projector $\hat{Q}_{x,x+1}$ is unitary equivalent to Eqs.
+Figure 1: The illustration of the 2D model and its decomposition into 1D models.
+The seminal work by Bravyi and Gosset Bravyi and Gosset (2015) showed that the Hamiltonian $\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-1}\hat{Q}_{x,x+1}$ under open boundary condition (OBC) can be gapless only when the projector $\hat{Q}_{x,x+1}$ is unitary equivalent to Eqs.
 (7), (8) for the rank 1 case and Eq.
 (13) for the rank 2 case.
-Then the min-max principle implies the same for $\hat{H}^{\mathrm{pbc}}$ at least when $\tilde{N}_{\mathrm{deg}}$ under OBC is finite, since $\hat{H}^{\mathrm{pbc}}-\hat{H}^{\mathrm{obc}}=\hat{Q}_{L,1}\geq 0$ is positive semi-definite.
-We will see that the fully polarized state
-$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x=1}^{L}|0\rangle_{x}=|0\cdots 0\rangle$ (15)
-is a common ground state of these Hamiltonians. Furthermore, the plane-wave state of a single spin flip
-$|{\Psi_{k}}\rangle\coloneqq\frac{1}{\sqrt{L}}\sum_{x=1}^{L}e^{ikx}|x\rangle,$ (16)
-$|x\rangle\coloneqq\hat{s}_{x}^{-}|{\Phi_{0}}\rangle=|\underbrace{0\cdots 0}_{x-1}1\underbrace{0\cdots 0}_{L-x}\rangle$ (17)
-with $k\coloneqq 2\pi m/L$ ($m=1,2,\cdots L$) is a common variational state among these examples and has a quadratic dispersion relation. That is,
-$E(k)$ $=\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\sum_{x=1}^{L}\big{\|}\hat{Q}_{x,x+1}|{\Psi_{k}}\rangle\big{\|}^{2}$
-$=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik(x-x^{\prime})}\langle x^{\prime}|\hat{H}^{\mathrm{pbc}}|x\rangle=O\big{(}(k-k_{0})^{2}\big{)}$ (18)
-for some $k_{0}$.
+Then the min-max principle implies the same for $\hat{H}$ under PBC, since $\hat{H}-\hat{H}^{\mathrm{obc}}=\hat{Q}_{L,1}\geq 0$ is positive semi-definite.
+We will see that the fully polarized state $|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x=1}^{L}|0\rangle_{x}=|0\cdots 0\rangle$ is a common ground state of these Hamiltonians.
+Furthermore, the plane-wave state
+$|{\Psi_{k}}\rangle\coloneqq\frac{1}{\sqrt{L}}\sum_{x=1}^{L}e^{ikx}\hat{s}_{x}^{-}|{\Phi_{0}}\rangle$ (6)
+with $k\coloneqq 2\pi m/L$ ($m=0,1,2,\cdots L-1$) is a common variational state with a quadratic dispersion relation.
 III.1 Rank 1 case
@@
 $\hat{Q}_{x,x+1}\coloneqq|\psi\rangle_{x,x+1}\langle\psi|_{x,x+1},$ (7)
-where $|\psi\rangle_{x,x+1}$ is a normalized state for two spins at $x$ and $x+1$. Without loss of generality one can assume the form
-$|\psi\rangle=(\alpha+i\beta)|01\rangle+(\alpha+i\gamma)|10\rangle+\delta|11\rangle,$ (20)
-$\alpha,\beta,\gamma,\delta\in\mathbb{R},\quad 2\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}=1$ (21)
-by performing a local unitary transformation (see Appendix B for the derivation) Bravyi and Gosset (2015). The model has U(1) spin rotation symmetry about the $z$ axis if and only if $\delta=0$.
-Since the matrix elements are known, the Hamiltonian in Eq.
-(19) can be expressed in terms of spin operators $\hat{s}_{x}^{a}$ and $\hat{s}_{x+1}^{a}$ ($a=x,y,z$).
-For example, when $(\alpha,\beta,\gamma,\delta)=\pm(0,1/\sqrt{2},-1/\sqrt{2},0)$, $\hat{Q}_{x,x+1}$ becomes the projector onto the spin-singlet state $(|01\rangle-|10\rangle)/\sqrt{2}$:
-$\hat{Q}_{x,x+1}$ $=\hat{\Pi}_{x,x+1}^{s=0},$ (22)
-$\hat{\Pi}_{x,x+1}^{s=0}$ $\coloneqq\hat{\mathbbm{1}}-\frac{1}{2}(\hat{\bm{s}}_{x}+\hat{\bm{s}}_{x+1})^{2}=\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (23)
-The corresponding Hamiltonian is the ferromagnetic Heisenberg model.
+where $|\psi\rangle_{x,x+1}$ is a normalized state for two spins at $x$ and $x+1$. Without loss of generality one can assume Bravyi and Gosset (2015)
+$|\psi\rangle=(\alpha+i\beta)|01\rangle+(\alpha+i\gamma)|10\rangle+\delta|11\rangle$ (8)
+with $\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $12\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}=$ 1, by performing a local unitary transformation (see Appendix B).
+For example, when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$, $\hat{Q}_{x,x+1}=\hat{\Pi}_{x,x+1}^{s=0}\coloneqq\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}$ is the projector onto the spin-singlet state $|s\rangle\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ and the corresponding Hamiltonian is the ferromagnetic Heisenberg model.
 Any state that is invariant under permutations of any two qubits (e.g., the fully-polarized state and the W state) is a ground state of this model.
-As a basis of such states, we can choose
-$\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L),$ (24)
-where $\hat{\mathcal{S}}$ is the symmetrization operator, uniformly averaging over the $\binom{L}{n}=\frac{L!}{(L-n)!n!}$ states.
-This model is known to be gapless due to the presence of a spin-wave excitation $E(k)=2\sin^{2}(k/2)$ under PBC.
-To analyze more general cases, let us introduce a matrix
-$m_{\psi}\coloneqq\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}=\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}.$ (25)
+As a basis of such states, we can choose $\{\hat{\mathcal{S}}|n\rangle\}_{n=0}^{L}$, where $|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle$ and $\hat{\mathcal{S}}$ is the symmetrization operator, uniformly averaging over the $\binom{L}{n}=\frac{L!}{(L-n)!n!}$ states.
+To analyze more general cases, let us introduce a matrix $m_{\psi}\coloneqq\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}$.
 The model defined by Eqs.
-(14), (19), and (20) can be gapless only when $\det m_{\psi}=(\alpha-i\beta)(-\alpha+i\gamma)\neq 0$, i.e., $|\psi\rangle$ is entangled Bravyi and Gosset (2015).
+(7), (8) can be gapless only when $\det m_{\psi}=(\alpha-i\beta)(-\alpha+i\gamma)\neq 0$, i.e., $|\psi\rangle$ is entangled Bravyi and Gosset (2015).
 In this case, we introduce an invertible operator $\hat{M}$ composed of nonuniform powers of the matrix $m_{\psi}$:
@@
 $\hat{m}_{\psi}\coloneqq(\alpha-i\beta)|0\rangle\langle 0|-(\alpha-i\gamma)|1\rangle\langle 1|+\delta|0\rangle\langle 1|.$ (10)
-This operator maps $|\psi\rangle$ to the singlet state:
-$\hat{M}^{\dagger}|\psi\rangle_{x,x+1}=(\det{m_{\psi}}^{*})^{j}(|01\rangle_{x,x+1}-|10\rangle_{x,x+1}),$ (28)
-It follows that $\hat{Q}_{x,x+1}$ for the state in Eq. (20) can be connected to $\hat{\Pi}_{x,x+1}^{s=0}$ as
+This operator maps $|\psi\rangle$ to the singlet state $\hat{M}^{\dagger}|\psi\rangle=(\det{m_{\psi}}^{*})^{j}|s\rangle$.
+It follows that $\hat{Q}_{x,x+1}$ for the state in Eq.
+(8) can be connected to $\hat{\Pi}_{x,x+1}^{s=0}$ as
 $\hat{Q}_{x,x+1}=4|{\det{m}_{\psi}}|^{2}(\hat{M}\hat{\Pi}_{x,x+1}^{s=0}\hat{M}^{-1})^{\dagger}(\hat{M}\hat{\Pi}_{x,x+1}^{s=0}\hat{M}^{-1}).$ (11)
@@
 (2021).
-As a consequence, ground states of $\hat{H}^{\mathrm{obc}}$ takes the form Bravyi and Gosset (2015)
-$\hat{M}\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L).$ (30)
+As a consequence, ground states of $\hat{H}^{\mathrm{obc}}$ takes the form $\hat{M}\hat{\mathcal{S}}|n\rangle$ ($n=0,1,\cdots L$Bravyi and Gosset (2015).
 The Hamiltonian under PBC has an additional term $\hat{Q}_{L,1}$.
-Every ground state of $\hat{H}^{\mathrm{obc}}$ remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if $m_{\psi}^{L}$ is proportional to the identity matrix $\mathbbm{1}$.
+Every ground state of $\hat{H}^{\mathrm{obc}}$ remains a ground state of $\hat{H}$ if $m_{\psi}^{L}$ is proportional to the identity matrix $\mathbbm{1}$.
 However, except when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$ discussed above, $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$ for some $L$.
-In such a case, the ground state of $\hat{H}^{\mathrm{pbc}}$ is restricted to the product state of the eigenstates of $\hat{m}_{\psi}$; that is, $|{\Phi_{0}}\rangle=|00\cdots 0\rangle$ and $|uu\cdots u\rangle$, where
-$|u\rangle\coloneqq\frac{[2\alpha-i(\beta+\gamma)]|1\rangle-\delta|0\rangle}{\sqrt{4\alpha^{2}+(\beta+\gamma)^{2}+\delta^{2}}}.$ (31)
-One can easily check that $\langle\psi|00\rangle=\langle\psi|uu\rangle=0$. Therefore, the exact ground state degeneracy of this model can be summarized as Bravyi and Gosset (2015)
-$\tilde{N}_{\mathrm{deg}}=\begin{cases}L+1&\text{if }m_{\psi}^{L}\propto\mathbbm{1},\\ 2&\text{otherwise}.\\ \end{cases}$ (32)
+In such a case, the ground state of $\hat{H}$ is restricted to the product state of the eigenstates of $\hat{m}_{\psi}$; that is, $|{\Phi_{0}}\rangle=|00\cdots 0\rangle$ and $|uu\cdots u\rangle$, where $|u\rangle\coloneqq\frac{[2\alpha-i(\beta+\gamma)]|1\rangle-\delta|0\rangle}{\sqrt{4\alpha^{2}+(\beta+\gamma)^{2}+\delta^{2}}}$.
 The variational state $|{\Psi_{k}}\rangle$ in Eq.
 (6) with $k\neq 0$ is orthogonal to the ground states at least when $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$, since their translation eigenvalues are different.
-The energy expectation value in Eq.
-(18) can be computed as
-$E(k)$ $=\sum_{x=1}^{L}\big{|}\langle\psi|_{x,x+1}|{\Psi_{k}}\rangle\big{|}^{2}=\big{|}\langle\psi|10\rangle+e^{ik}\langle\psi|01\rangle\big{|}^{2}$
-$=\big{|}(\alpha-i\gamma)+(\alpha-i\beta)e^{ik}\big{|}^{2}$
-$=(1-\delta^{2})+2(\alpha^{2}+\beta\gamma)\cos k+2\alpha(\beta-\gamma)\sin k$
-$=(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}\cos(k-k_{0}).$ (33)
-In going to the last line, we defined
-$\cos k_{0}\coloneqq-\frac{\alpha^{2}+\gamma\beta}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}},$ (34)
-$\sin k_{0}\coloneqq-\frac{\alpha(\beta-\gamma)}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}}.$ (35)
-It follows that $E(k)$ can be bounded below as
-$E(k)\geq(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}.$ (36)
-The right-hand side vanishes when $\beta^{2}=\gamma^{2}$, implying the existence of gapless excitations when $\gamma=\pm\beta$. In this case $E(k)$ can be rewritten as
-$E(k)=2(1-\delta^{2})\sin^{2}((k-k_{0})/2),$ (37)
-which confirms our conjecture in the rank 1 case.
+The energy expectation value $\langle\Psi_{k}|\hat{H}|\Psi_{k}\rangle$ is given by
+$E(k)$ $=\big{|}(\alpha-i\gamma)+(\alpha-i\beta)e^{ik}\big{|}^{2}$
+$=(1-\delta^{2})-A\cos(k-k_{0}),$ (12)
+where $A\coloneqq\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}$ and $k_{0}$ is defined by $\cos k_{0}=-2(\alpha^{2}+\gamma\beta)/A$ and $\sin k_{0}=-2\alpha(\beta-\gamma)/A$.
+This excitation energy takes the minimum value at $k=k_{0}$, which vanishes if $\beta^{2}=\gamma^{2}$, implying the existence of gapless excitations.
+In this case $E(k)$ can be rewritten as $E(k)=2(1-\delta^{2})\sin^{2}((k-k_{0})/2)$, confirming our conjecture in this model.
 When $\gamma=\beta$, $k_{0}$ is precisely $\pi$, but it takes more general value when $\gamma=-\beta$.
-This result is consistent with Ref.
-Bravyi and Gosset (2015) which mathematically proved that $\hat{H}^{\mathrm{obc}}$ is gapped if and only if $\beta^{2}\neq\gamma^{2}$.
 III.2 Rank 2 case
-Next let us discuss the rank 2 case. In this case the Hamiltonian in Eq. (14) can be gapless only when $\hat{Q}_{x,x+1}$ is unitary equivalent to the XY model with an external magnetic field:
-$\hat{Q}_{x,x+1}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\frac{\zeta\hat{s}_{x}^{+}\hat{s}_{x+1}^{-}+\zeta^{*}\hat{s}_{x}^{-}\hat{s}_{x+1}^{+}}{1+|\zeta|^{2}}-\frac{|\zeta|^{2}\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}}{{1+|\zeta|^{2}}}.$ (38)
-with $\zeta\in\mathbb{C}$ Bravyi and Gosset (2015). One may express this Hamiltonian as
-$\mathrm{ker}\,\hat{Q}_{x,x+1}=\mathrm{span}\big{\{}|00\rangle,|10\rangle+\zeta|01\rangle\big{\}}.$ (39)
+Next we discuss the rank 2 case. This time the Hamiltonian can be gapless only when $\hat{Q}_{x,x+1}$ is unitary equivalent to the XY model with an external magnetic field Bravyi and Gosset (2015):
+$\hat{Q}_{x,x+1}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\frac{\zeta\hat{s}_{x}^{+}\hat{s}_{x+1}^{-}+\text{h.c.}}{1+|\zeta|^{2}}-\frac{|\zeta|^{2}\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}}{{1+|\zeta|^{2}}}$ (13)
+with $\zeta\in\mathbb{C}$.
+This Hamiltonian can also be expressed as $\mathrm{ker}\,\hat{Q}_{x,x+1}=\mathrm{span}\big{\{}|00\rangle,|10\rangle+\zeta|01\rangle\big{\}}$.
 When $\zeta=1$, ground states of $\hat{Q}_{x,x+1}^{\zeta=1}$ are the fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of single spin flip $|\Psi_{0}\rangle$, regardless of the boundary condition.
 We find that $\hat{Q}_{x,x+1}$ with $\zeta\neq 1$ can be obtained from $\hat{Q}_{x,x+1}^{\zeta=1}$ by Witten’s conjugation as
-$\hat{Q}_{x,x+1}=(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1})^{\dagger}\hat{C}_{x,x+1}(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1}),$ (40)
-where
-$\hat{M}\coloneqq\hat{\mathbbm{1}}\otimes\hat{m}_{\zeta}\otimes\hat{m}_{\zeta}^{2}\otimes\cdots\otimes\hat{m}_{\zeta}^{L-1},$ (41)
-$\hat{m}_{\zeta}\coloneqq|0\rangle\langle 0|+\zeta|1\rangle\langle 1|,$ (42)
-$\hat{C}\coloneqq\hat{\mathbbm{1}}+\frac{(1-|\zeta|^{2})^{2}}{2\mathrm{Re}\zeta(1+|\zeta|^{2})}(|10\rangle\langle 01|+|01\rangle\langle 10|).$ (43)
-Hence, the ground states of $\hat{H}^{\mathrm{obc}}$ with $\zeta\neq 1$ are given by $\hat{M}|{\Phi_{0}}\rangle=|{\Phi_{0}}\rangle$ and $\hat{M}|\Psi_{0}\rangle\propto\sum_{x=1}^{L}\zeta^{x-1}|x\rangle.$ The latter remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if and only if $\zeta^{L}=1$.
-Because of the U(1) symmetry of the Hamiltonian in Eq.
-(38), the plane-wave state $|{\Psi_{k}}\rangle$ in Eq.
-(16) is actually an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$.
-The eigen energy is given by
-$E(k)$ $=\frac{|e^{ik}-\zeta|^{2}}{1+|\zeta|^{2}}=\frac{\big{|}e^{i(k-k_{0})}-|\zeta|\big{|}^{2}}{1+|\zeta|^{2}}$
-$=\frac{1-|\zeta|^{2}}{1+|\zeta|^{2}}+\frac{4|\zeta|}{1+|\zeta|^{2}}\sin^{2}((k-k_{0})/2),$ (44)
-where $k_{0}$ is defined by $\zeta=|\zeta|e^{ik_{0}}$. Hence, the model is gapless and the conjecture holds when $|\zeta|=1$, regardless of $\zeta^{L}=1$ or not.
-III.3 Degeneracy in gapped cases
-Although our main focus is on gapless models, let us discuss the ground state degeneracy of gapped cases for the model in Eq.
-(14).
-This result will be used in our discussion for higher dimensional models in Sec.
-IV.
-Let us start with the rank 1 model.
-In the $\gamma^{2}\neq\beta^{2}$ case, the formula Eq.
-(32) implies that $N_{\mathrm{deg}}=2$, because $m_{\psi}^{L}$ cannot be proportional to $\mathbbm{1}$ when the absolute values of the diagonal elements $\alpha-i\beta$ and $-\alpha+i\gamma$ do not agree.
-The $\alpha=\beta=\gamma=0$ case is exceptional.
-In this case $|\psi\rangle=\pm|11\rangle$ is a symmetric product state and any product state with no consecutive $1$’s is a ground state.
-Hence, the ground state degeneracy under PBC with $L\geq 3$ is given by the Lucas number Wikipedia article
-$N_{\mathrm{deg}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2},$ (45)
-which increases exponentially with the system size $L$. Here $\lfloor x\rfloor$ represents the greatest integer less than or equal to $x\in\mathbb{R}$. We append the derivation in Appendix C.
-Next, let us discuss the rank 2 case.
-When $|\zeta|\neq 1$, the ground state of $\hat{H}^{\mathrm{pbc}}$ composed of the projector in Eq.
-(38) is unique ($N_{\mathrm{deg}}=1$) because $\zeta^{L}$ cannot be 1.
-According to Ref.
-Bravyi and Gosset (2015), $N_{\mathrm{deg}}\leq 2$ holds even for other rank 2 models under OBC.
-The min-max principle in Eq.
-(12) implies the same under PBC.
-In summary, for translation-invariant $s=1/2$ chains with nearest neighbor interactions, the ground state degeneracy in the gapped cases is bounded as
-$N_{\mathrm{deg}}\leq 2,$ (46)
-except for the rank 1 model with $\alpha=\beta=\gamma=0$.
-Figure 1: The illustration of the 2D model in Eq. (47) and its decomposition into 1D models.
+$\hat{Q}_{x,x+1}=(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1})^{\dagger}\hat{C}(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1}),$ (14)
+where $\hat{M}\coloneqq\hat{\mathbbm{1}}\otimes\hat{m}_{\zeta}\otimes\cdots\otimes\hat{m}_{\zeta}^{L-1}$, $\hat{m}_{\zeta}\coloneqq|0\rangle\langle 0|+\zeta|1\rangle\langle 1|$, and $\hat{C}\coloneqq\hat{\mathbbm{1}}+\frac{(1-|\zeta|^{2})^{2}}{2\mathrm{Re}\zeta(1+|\zeta|^{2})}(|10\rangle\langle 01|+|01\rangle\langle 10|)$.
+Hence, the ground states of $\hat{H}^{\mathrm{obc}}$ with $\zeta\neq 1$ are given by $\hat{M}|{\Phi_{0}}\rangle=|{\Phi_{0}}\rangle$ and $\hat{M}|\Psi_{0}\rangle\propto\sum_{x=1}^{L}\zeta^{x-1}|x\rangle.$ The latter remains a ground state under PBC if and only if $\zeta^{L}=1$.
+Because of the U(1) symmetry of the Hamiltonian in Eq. (13), the plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (6) is actually an exact eigenstate of $\hat{H}$. The eigenenergy is
+$E(k)=\frac{|e^{ik}-\zeta|^{2}}{1+|\zeta|^{2}}=\frac{1-|\zeta|^{2}}{1+|\zeta|^{2}}+\frac{4|\zeta|}{1+|\zeta|^{2}}\sin^{2}\Big{(}\frac{k-k_{0}}{2}\Big{)},$ (15)
+where $k_{0}$ is defined by $\zeta=|\zeta|e^{ik_{0}}$. Hence, the model is gapless and the conjecture holds when $|\zeta|=1$.
+Finally, for the gapped cases, we found that the ground state degeneracy is, in general, bounded as $N_{\mathrm{deg}}\leq 2$, except the rank 1 model with $\alpha=\beta=\gamma=0$ for which
+$N_{\mathrm{deg}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2}.$ (16)
+See Appendix C for the derivation. This coincides with the Lucas number Wikipedia article that increases exponentially with the system size $L$.
 IV 2D models with nearest-neighbor interactions
-In this section, we generalize the results in the previous section to two and three dimensional models using the min-max principle.
-We consider a square lattice of $s=1/2$ spins and assume the Hamiltonian of the form:
-$\hat{H}^{\mathrm{2D}}\coloneqq\sum_{x,y=1}^{L}\big{(}\hat{Q}_{(x,y),(x+1,y)}+\hat{Q}_{(x,y),(x,y+1)}\big{)},$ (47)
-where $\hat{Q}_{(x,y),(x^{\prime},y^{\prime})}$ is a projector acting on two spins at $(x,y)$ and $(x^{\prime},y^{\prime})$ [Fig.
+In this section, we generalize results obtained in Sec.
+III to higher dimensions.
+We consider a square lattice of $s=1/2$ spins, assuming the Hamiltonian of the form $\hat{H}\coloneqq\sum_{x,y=1}^{L}\big{(}\hat{Q}_{(x,y),(x+1,y)}+\hat{Q}_{(x,y),(x,y+1)}\big{)}$, where $\hat{Q}_{(x,y),(x^{\prime},y^{\prime})}$ is a projector acting on two spins at $(x,y)$ and $(x^{\prime},y^{\prime})$ [Fig.
 1 (a)].
 We impose PBC and assume the translation invariance in both $x$ and $y$ directions.
-Our goal is to show that when $\hat{H}^{\mathrm{2D}}$ is frustration free and gapless, there exists a variational state with properties in Eqs.
-(8) and (9).
-Let us decompose the 2D Hamiltonian into 1D chains in two different ways.
-$\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}.$ (48)
-The first choice is the simple one, decomposing the system into decoupled chains as illustrated in Fig. 1 (b) and (c):
-$\hat{H}^{(1)}\coloneqq\sum_{y=1}^{L}\hat{H}_{y}^{\mathrm{1D}},\quad\hat{H}_{y}^{\mathrm{1D}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{(x,y),(x+1,y)},$ (49)
-$\hat{H}^{(2)}\coloneqq\sum_{x=1}^{L}\hat{H}_{x}^{\mathrm{1D}},\quad\hat{H}_{x}^{\mathrm{1D}}\coloneqq\sum_{y=1}^{L}\hat{Q}_{(x,y),(x,y+1)}.$ (50)
-For each $y$, $\hat{H}_{y}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $x$ axis; similarly, for each $x$, $\hat{H}_{x}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $y$ axis.
-The second choice is illustrated in Fig.
-1 (d) and (e), in which the chains are connected into one piece.
-Even when these 1D Hamiltonians are frustration-free, whether the total Hamiltonian $\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}$ remains frustration-free or not is, in general, nontrivial.
+Our goal is to show that, if $\hat{H}$ is FF and gapless, there exists a variational state $|\Psi_{\bm{k}}\rangle$ with the dispersion satisfying (4).
+To this end, let us decompose the 2D Hamiltonian into 1D chains in two different ways.
+The first choice is the simple one, decomposing $\hat{H}$ into decoupled chains $\hat{H}^{(1)}\coloneqq\sum_{y=1}^{L}\hat{H}_{y}^{\mathrm{1D}}$ and $\hat{H}^{(2)}\coloneqq\sum_{x=1}^{L}\hat{H}_{x}^{\mathrm{1D}}$ as illustrated in Fig.
+1 (b,c).
+For each $y$, $\hat{H}_{y}^{\mathrm{1D}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{(x,y),(x+1,y)}$ describes the Hamiltonian for the chain along the $x$ axis; similarly, for each $x$, $\hat{H}_{x}^{\mathrm{1D}}\coloneqq\sum_{y=1}^{L}\hat{Q}_{(x,y),(x,y+1)}$ describes the Hamiltonian for the chain along the $y$ axis.
+The second decomposition $\hat{H}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}$ is illustrated in Fig.
+1 (d,e), in which the chains are connected into one piece.
+Even when these 1D Hamiltonians are FF, whether the total Hamiltonian $\hat{H}$ remains FF or not is, in general, nontrivial.
 However, this issue can be easily gone around in our case.
-When either $\hat{H}^{\prime}{}^{(1)}$ or $\hat{H}^{\prime}{}^{(2)}$ is gapped with a finite ground state degeneracy, the min-max principle suggests that $\hat{H}^{\mathrm{2D}}$ is also gapped, because $\hat{H}^{\prime}{}^{(2)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(1)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(2)}$ are positive semi-definite  [^1] This argument cannot be directly applied to the simpler decomposition $\hat{H}^{(1)}+\hat{H}^{(2)}$, because if $N_{\mathrm{deg}}\geq 2$ for each $\hat{H}_{y}^{\mathrm{1D}}$, then the ground state degeneracy of $\hat{H}^{(1)}$ becomes $N_{\mathrm{deg}}^{L}$, which grows exponentially with the system size and the min-max principle becomes silent.
+When either $\hat{H}^{\prime}{}^{(1)}$ or $\hat{H}^{\prime}{}^{(2)}$ is gapped with a finite ground state degeneracy, the min-max principle suggests that $\hat{H}$ is also gapped, because $\hat{H}^{\prime}{}^{(2)}=\hat{H}-\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(1)}=\hat{H}-\hat{H}^{\prime}{}^{(2)}$ are positive semi-definite  [^1] This argument cannot be directly applied to the simpler decomposition $\hat{H}^{(1)}+\hat{H}^{(2)}$, because if $N_{\mathrm{deg}}\geq 2$ for each $\hat{H}_{y}^{\mathrm{1D}}$, then the ground state degeneracy of $\hat{H}^{(1)}$ becomes $N_{\mathrm{deg}}^{L}$, which grows exponentially with the system size and the min-max principle becomes silent.
 For the same reason, the case of $\alpha=\beta=\gamma=0$ in the rank 1 model should be treated separately, but this case is trivially gapped in any spatial dimension.
@@
 Hence, we can restrict ourselves to the case where both $\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(2)}$ are gapless, implying that $\hat{H}^{(1)}$ and $\hat{H}^{(2)}$ are also gapless.
-Furthermore, as we have seen in Sec.
-III, when the Hamiltonian for an $s=1/2$ spin chain with nearest-neighbor interaction is gapless, the fully polarized state $|{\Phi_{0}}\rangle$ in Eq.
-(15) is a common zero-energy ground state among all models.
-Hence, the 2D version of the fully polarized state
-$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x,y=1}^{L}|0\rangle_{(x,y)}$ (51)
-is a simultaneous ground state of all terms in $\hat{H}^{\mathrm{2D}}$, implying that $\hat{H}^{\mathrm{2D}}$ is frustration-free.
-To construct a variational state for low-energy excitations, let us define the plane-wave state of a single spin flip by
-$|\Phi_{(k_{x},k_{y})}\rangle\coloneqq\frac{1}{L}\sum_{x,y=1}^{L}e^{i(k_{x}x+k_{y}y)}\hat{s}_{(x,y)}^{-}|{\Phi_{0}}\rangle.$ (52)
+Furthermore, as we have seen above, when the Hamiltonian for an $s=1/2$ spin chain with nearest-neighbor interaction is gapless, the fully polarized state $|{\Phi_{0}}\rangle$ is a common zero-energy ground state.
+Hence, the 2D version of the fully polarized state $|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x,y=1}^{L}|0\rangle_{(x,y)}$ is a simultaneous ground state of all terms in $\hat{H}$, implying that $\hat{H}$ is FF.
+To construct a variational state for low-energy excitations, let us define the plane-wave state of a single spin flip by $|\Psi_{\bm{k}}\rangle\coloneqq\frac{1}{L}\sum_{x,y=1}^{L}e^{i\bm{k}\cdot\bm{r}}\hat{s}_{(x,y)}^{-}|{\Phi_{0}}\rangle$.
 The variational energy is given by
-$E(k_{x},k_{y})$ $\coloneqq\langle\Phi_{(k_{x},k_{y})}|\hat{H}^{\mathrm{2D}}|\Phi_{(k_{x},k_{y})}\rangle$
-$=\sum_{y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{y}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
-$\quad+\sum_{x=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{x}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
-$=E(k_{x})+E(k_{y}),$ (53)
-where we used the relation
-$E(k_{x})=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik_{x}(x-x^{\prime})}\langle x^{\prime},y|\hat{H}_{y}^{\mathrm{1D}}|x,y\rangle,$ (54)
-$E(k_{y})=\frac{1}{L}\sum_{y,y^{\prime}=1}^{L}e^{ik_{y}(y-y^{\prime})}\langle x,y^{\prime}|\hat{H}_{x}^{\mathrm{1D}}|x,y\rangle.$ (55)
-Therefore,
-$E(k_{x},k_{y})=O\big{(}(k_{x}-k_{0})^{2},(k_{y}-k_{0})^{2}\big{)}$ (56)
-The discussion above can be readily extended to the cubic lattice in three dimensions.
-The variational energy of the plane-wave state $|\Phi_{(k_{x},k_{y},k_{z})}\rangle\coloneqq L^{-3/2}\sum_{x,y,z=1}^{L}e^{i(k_{x}x+k_{y}y+k_{z}z)}\hat{s}_{(x,y,z)}^{-}|{\Phi_{0}}\rangle$ is simply given by $E(k_{x})+E(k_{y})+E(k_{z})$.
-Finally, let us discuss whether this analysis can be generalized to other form of lattices. For example, the nearest-neighbor interaction on the triangular lattice contains a diagonal interaction
-$\sum_{x,y=1}^{L}\hat{Q}_{(x,y),(x+1,y-1)}.$ (57)
-More generally, when an interaction $\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}$ among the spins at $(x,y)$ and $(x+d_{x},y+d_{y})$ is added, the variational energy obtains a term
-$E(k_{x}d_{x}+k_{y}d_{y})$
-$=\sum_{x,y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}|\Phi_{(k_{x},k_{y})}\rangle.$ (58)
-Hence, the variational state generically becomes gapped unless $E(k_{0}(d_{x}+d_{y}))=0$.
-When $k_{0}=0$ (for example, in the ferromagnetic Heisenberg model), this condition can be easily satisfied, but otherwise some fine-tuning is required.
-V 1D examples with next nearest-neighbor interactions
-In Sec.
-III, we showed that frustration-free Hamiltonians for $s=1/2$ spin chains with nearest-neighbor interactions have a common ground state and a low-energy variational state.
-This observation was the key in demonstrating the softness of the dispersion relation in higher dimensional models in Sec.
-IV.
-Table 1: Comparison of gapless frustration-free Hamiltonians for $s=1/2$ chain.
-$|{\Phi_{0}}\rangle$ [(15)] is the fully polarized state and $|{\Psi_{k}}\rangle$ [Eq.
+$E(\bm{k})$ $\coloneqq\langle\Psi_{\bm{k}}|\hat{H}|\Psi_{\bm{k}}\rangle=E(k_{x})+E(k_{y}),$ (17)
+which satisfies (4).
+This discussion can be readily extended to the cubic lattice.
+The variational energy of the plane-wave state $|\Psi_{\bm{k}}\rangle\coloneqq L^{-3/2}\sum_{x,y,z=1}^{L}e^{i\bm{k}\cdot\bm{r}}\hat{s}_{(x,y,z)}^{-}|{\Phi_{0}}\rangle$ is simply given by $E(k_{x})+E(k_{y})+E(k_{z})$.
+Finally, let us investigate possible generalization to other form of lattices.
+For example, the nearest-neighbor interaction on the triangular lattice contains a diagonal interaction $\sum_{x,y=1}^{L}\hat{Q}_{(x,y),(x+1,y-1)}$.
+More generally, when an interaction $\hat{Q}_{(x,y),(x+d_{x},y+d_{y})}$ among the spins at $(x,y)$ and $(x+d_{x},y+d_{y})$ is added, the variational energy obtains a term $E(k_{x}d_{x}+k_{y}d_{y})$.
+Hence, the plane-wave state generically becomes gapped unless $E(k_{0}(d_{x}+d_{y}))=0$.
+When $k_{0}=0$ (e.g., in the ferromagnetic Heisenberg model), this condition can be easily satisfied, but otherwise some fine-tuning is required.
+V Examples
+In this section we discuss several examples for which our general discussing in Sec. III and IV are not applicable. We summarize our results in Table 1.
+Table 1: Comparison of gapless FF Hamiltonians for $s=1/2$ chain.
+$|{\Phi_{0}}\rangle=|00\cdots\rangle$ is the fully polarized state and $|{\Psi_{k}}\rangle$ [Eq.
 (6)] is the plane-wave state of single spin flip.
+The XYZ MG model is discussed in Sec.
+IV of SM.
 Interaction type Range $|{\Phi_{0}}\rangle$ is a GS $|{\Psi_{k}}\rangle$ is gapless Variational energy
@@
 Nearest neighbor models with rank 2: (13) with $|\zeta|=1$ 1 $O\big{(}(k-k_{0})^{2}\big{)}$
-Fredkin spin chain: (63) 2 $O\big{(}k^{2}\big{)}$
 GHZ uncle Hamiltonian: (19) 2 $O\big{(}k^{2},L^{-2}\big{)}$
-XYZ MG model: (78), (82) with $\theta\neq 0$ 2 $O\big{(}k^{2})$
-In this section, we discuss $s=1/2$ spin chains with longer-range interactions. We assume that the Hamiltonian has the following form
-$\hat{H}^{\mathrm{pbc}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{x,x+1,x+2},\quad\hat{H}^{\mathrm{obc}}\coloneqq\sum_{x=1}^{L-2}\hat{Q}_{x,x+1,x+2}.$ (59)
-Here, $\hat{Q}_{x,x+1,x+2}$ is a projector acting on the three spins at $x$, $x+1$, and $x+2$.
-The interaction range of the Hamiltonian is 2.
-Unlike the only nearest-neighbor interaction case in the previous section, our discussion here is based on examples and hence is not comprehensive.
-However, these examples clearly demonstrate difficulties in the direct extension of our results to more general cases.
-Namely, the fully polarized state $|{\Phi_{0}}\rangle$ may not be a ground state and the plane-wave state $|\Psi_{k}\rangle$ may not be a low-energy state.
-We instead construct a low-energy variational state which does not take the form $\sum_{x}e^{ikx}\hat{O}_{x}|\Phi_{0}\rangle$.
-We summarize our results in Table 1.
-V.1 W state
-Let us start with the uncle Hamiltonian constructed by the fully polarized state and the W state $|W\rangle\coloneqq(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ Fernández-González et al.
+XYZ MG model with $\theta\neq 0$: (29) 2 $O\big{(}k^{2})$
+V.1 Kinetic Ising model
+Let us consider an $s=1/2$ spin model defined on a $d$-dimensional cubic lattice: $\hat{H}=\sum_{\bm{r}}\hat{H}_{\bm{r}}$, where
+$\hat{H}_{\bm{r}}=\frac{1}{2\cosh(J\sum_{\bm{r}^{\prime}\in B_{\bm{r}^{\prime}}}\hat{\sigma}_{\bm{r}^{\prime}}^{z})}\Big{(}e^{-J\hat{\sigma}_{\bm{r}}^{z}{}\sum_{\bm{r}^{\prime}\in B_{\bm{r}}}\hat{\sigma}_{\bm{r}^{\prime}}^{z}}-\hat{\sigma}_{\bm{r}}^{x}\Big{)}.$ (18)
+Here, $\hat{\sigma}_{\bm{r}}^{a}$ ($a=x,y,z$) are the Pauli matrices related to spin operators by $\hat{s}_{\bm{r}}^{a}=(1/2)\hat{\sigma}_{\bm{r}}^{a}$ and $B_{\bm{r}}$ is the nearest neighbors of $\bm{r}$.
+This model is associated with the classical Ising model with the Boltzmann weight $\propto e^{J\sum_{\langle\bm{r},\bm{r}^{\prime}\rangle}\sigma_{\bm{r}}^{z}\sigma_{\bm{r}^{\prime}}^{z}}$ Masaoka et al.
 (2024).
-The kernel of the Hamiltonian is
-$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}=\mathrm{span}\big{\{}|000\rangle,|W\rangle\big{\}}.$ (60)
-The spin representation of the Hamiltonian is found to be
-$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{3}{4}\hat{\mathbbm{1}}-\frac{\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}+\hat{s}_{x+2}^{z}}{3}+2\hat{s}_{x}^{z}\hat{s}_{x+1}^{z}\hat{s}_{x+2}^{z}$
-$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}$
-$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+1}^{z}\Big{)}\hat{\bm{s}}_{x+2}\cdot\hat{\bm{s}}_{x}$
-$\quad-\frac{2}{3}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x+2}^{z}\Big{)}\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (61)
-The fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of a single spin flip $|\Psi_{0}\rangle$ are the two ground states of this model regardless of the boundary condition.
-The state $|{\Psi_{k}}\rangle$ in Eq.
-(16) is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ as expected by the U(1) symmetry.
-The eigen energy is given by
-$E(k)$ $=\big{|}\hat{Q}_{1,2,3}(|100\rangle+e^{ik}|010\rangle+e^{2ik}|001\rangle)\big{|}^{2}$
-$=\big{|}\hat{Q}_{1,2,3}((e^{ik}-1)|010\rangle+(e^{2ik}-1)|001\rangle)\big{|}^{2}$
-$=2(1-\cos k)\big{|}\hat{Q}_{1,2,3}(|010\rangle+(e^{ik}+1)|001\rangle)\big{|}^{2}$
-$=\frac{8}{3}(2+\cos k)\sin^{2}(k/2).$ (62)
-As a related model, let us discuss the bulk Hamiltonian for the Fredkin spin chain Salberger and Korepin (2017):
-$\hat{Q}_{x,x+1,x+2}^{\prime}$ $\coloneqq\Big{(}\frac{1}{2}\hat{\mathbbm{1}}+\hat{s}_{x}^{z}\Big{)}\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}\Big{)}$
-$\quad+\Big{(}\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}\Big{)}\Big{(}\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+2}^{z}\Big{)}.$ (63)
-The kernel of this Hamiltonian is
-$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}$ $=\mathrm{span}\Big{\{}|000\rangle,|100\rangle,|001\rangle+|010\rangle,$
-[…]
-Since $\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}^{\prime}\supset\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$, the min-max principle implies that the Fredkin spin chain is gapless with dispersion equal to or softer than the one implied by Eq.
+(The temperature $T$ is included in the definition of $J$.) The $\mathbb{Z}_{2}$ symmetry $\prod_{\bm{r}}\hat{\sigma}_{\bm{r}}^{x}$ of the model is spontaneously broken only in the limit $J\to\infty$, which corresponds to the zero temperature limit, in one dimension and for $J$ above a critical value $J_{c}$ in higher dimensions.
+V.1.1 1D
+In one dimension, $\hat{H}_{x}$ is a projector $\hat{Q}_{x-1,x,x+1}$ acting on the three spins at $x-1$, $x$, and $x+1$, which can be expressed as
+$\hat{Q}_{x-1,x,x+1}$
+$=\frac{1}{2}\hat{\mathbbm{1}}-c_{2}\hat{s}_{x}^{z}(\hat{s}_{x-1}^{z}+\hat{s}_{x+1}^{z})-2\hat{s}_{x}^{x}\Big{(}\frac{c_{1}}{4}-c_{3}\hat{s}_{x-1}^{z}\hat{s}_{x+1}^{z}\Big{)}$ (19)
+with $c_{1}=\frac{2\cosh^{2}J}{\cosh(2J)}$, $c_{2}=\tanh(2J)$, and $c_{3}=\frac{2\sinh^{2}J}{\cosh(2J)}$.
+In the $J\to\infty$ limit, all three parameters $c_{1}$, $c_{2}$, $c_{3}$ become $1$ and the model reduces to the uncle Hamiltonian for the Greenberger–Horne–Zeilinger (GHZ) state Fernández-González et al.
 (2015).
-Indeed, $|{\Psi_{k}}\rangle$ is an exact eigenstate with eigenenergy
-$E^{\prime}(k)$ $=1-\cos k=2\sin^{2}(k/2).$ (65)
-In these two examples, the fully polarized state $|{\Phi_{0}}\rangle$ is one of the ground states and the plane-wave state $|\Psi_{k}\rangle$ is a low-energy state with the dispersion $E(k)=O(k^{2})$.
-V.2 Greenberger–Horne–Zeilinger uncle Hamiltonian
-Next, we discuss the uncle Hamiltonian for the Greenberger–Horne–Zeilinger (GHZ) state Fernández-González et al. (2015). The kernel of the Hamiltonian is
-$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
-$=\mathrm{span}\Big{\{}|000\rangle,|111\rangle,|100\rangle+|110\rangle,|001\rangle+|011\rangle\Big{\}}.$ (66)
-The spin representation the Hamiltonian is
-$\hat{Q}_{x,x+1,x+2}$ $\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\hat{s}_{x+1}^{z}(\hat{s}_{x}^{z}+\hat{s}_{x+2}^{z})$
-$\quad-2\hat{s}_{x+1}^{x}\Big{(}\frac{1}{4}-\hat{s}_{x}^{z}\hat{s}_{x+2}^{z}\Big{)}.$ (67)
-To write down ground states and low-energy excitations of this model, let us define
-$|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle,\quad|\bar{n}\rangle\coloneqq|\underbrace{1\cdots 1}_{L-n}\underbrace{0\cdots 0}_{n}\rangle.$ (68)
-for $n=1,2,\cdots,L-1$. For example, $|n\rangle$ satisfies
-$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
-$=-\frac{1}{2}(|n+1\rangle+|n-1\rangle+\hat{T}|n+1\rangle+\hat{T}^{-1}|n-1\rangle)$ (69)
+The kernel of $\hat{Q}_{x,x+1,x+2}$ is given by
+$\mathrm{span}\Big{\{}|000\rangle,|111\rangle,|100\rangle+|110\rangle,|001\rangle+|011\rangle\Big{\}}.$ (20)
+Figure 2: (a) The all spin-up state $|\Phi_{0}\rangle$ of 1D kinetic Ising model.
+(b) An excited state $|\Psi_{k,\ell}\rangle$ with two domain walls in Eq.
+(24).
+(c) A ground state of the 2D kinetic Ising model with straight domain walls.
+(d) One-dimensional gapless excitations on the surface of domain walls that corresponds to (b).
+To write down ground states and low-energy excitations of this model, let us introduce $|\bar{n}\rangle\coloneqq|\underbrace{1\cdots 1}_{L-n}\underbrace{0\cdots 0}_{n}\rangle$ in addition to $|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle$.
+For example, $|n\rangle$ satisfies
+$\hat{H}|n\rangle-2|n\rangle$
+$=-\frac{1}{2}(|n+1\rangle+|n-1\rangle+\hat{T}_{1}|n+1\rangle+\hat{T}_{1}^{-1}|n-1\rangle)$ (21)
 for $n=2,3,\cdots,L-2$, and
-$\hat{H}^{\mathrm{pbc}}|n\rangle-2|n\rangle$
-=$=\begin{cases}-\frac{1}{2}(|2\rangle+\hat{T}|2\rangle)&(n=1),\\ -\frac{1}{2}(|L-2\rangle+\hat{T}^{-1}|L-2\rangle)&(n=L-1)\end{cases}$ (70)
-for $n=1$ and $L-1$. Based on these expressions, the ground states of $\hat{H}^{\mathrm{obc}}$ are found to be
+$\hat{H}|n\rangle-2|n\rangle$
+=$=\begin{cases}-\frac{1}{2}(|2\rangle+\hat{T}_{1}|2\rangle)&(n=1),\\ -\frac{1}{2}(|L-2\rangle+\hat{T}_{1}^{-1}|L-2\rangle)&(n=L-1)\end{cases}$ (22)
+for $n=1$ and $L-1$. Based on these expressions, the ground states under OBC are found to be
 $|0\cdots 0\rangle\,(=|\Phi_{0}\rangle),|1\cdots 1\rangle,\sum_{n=1}^{L-1}|n\rangle,\sum_{n=1}^{L-1}|\bar{n}\rangle.$ (23)
-The latter two are not consistent with the PBC and are hence not ground states of $\hat{H}^{\mathrm{pbc}}$.
-Although the fully-polarized state $|{\Phi_{0}}\rangle$ is one of the ground states, the simple variational state $|{\Psi_{k}}\rangle$ is not a low-energy state:
-$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=2.$ (72)
-We instead consider a plane-wave state of two domain walls:
-$|\Psi_{k,\ell}\rangle\coloneqq\frac{\sqrt{2}}{L}\sum_{n=1}^{L-1}\sum_{m=0}^{L-1}e^{ikm}e^{-ikn/2}\sin(\tfrac{\pi n\ell}{L})\hat{T}^{m}|n\rangle$ (73)
-for $\ell=1,2,\cdots,L-1$.
-Using $\hat{T}|\Psi_{k,\ell}\rangle=e^{-ik}|\Psi_{k,\ell}\rangle$ and $[\hat{T},\hat{H}^{\mathrm{pbc}}]=0$, one can readily show that $|\Psi_{k,\ell}\rangle$ is an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$ with the eigenvalue
-$E_{\ell}(k)$ $=2[1-\cos(k/2)\cos(\tfrac{\pi\ell}{L})]$
-$=4\sin^{2}(\tfrac{\pi\ell}{2L})+4\cos(\tfrac{\pi\ell}{L})\sin^{2}(k/4).$ (74)
-The first term in the last expression is $O(L^{-2})$, which vanishes in the thermodynamic limit. The second term is $O(k^{2})$.
-Despite the presence of gapless excitations, in this model, the correlation function of the form
+The latter two are not consistent with the PBC and are hence not ground states of $\hat{H}$.
+Although the fully-polarized state $|{\Phi_{0}}\rangle$ [Fig 2(a)] is one of the ground states, the plane-wave state in Eq.
+(6) is not a low-energy state because $\langle\Psi_{k}|\hat{H}|{\Psi_{k}}\rangle=2$.
+We instead consider a plane-wave state of two domain walls illustrated in Fig 2(b):
+$|\Psi_{k,\ell}\rangle\coloneqq\frac{\sqrt{2}}{L}\sum_{n=1}^{L-1}\sum_{m=0}^{L-1}e^{ikm}e^{-ikn/2}\sin(\tfrac{\pi n\ell}{L})\hat{T}_{1}^{m}|n\rangle$ (24)
+for $\ell=1,2,\cdots,L-1$. We find that $|\Psi_{k,\ell}\rangle$ is an exact eigenstate of $\hat{H}$ and $\hat{T}_{1}$ with the energy eigenvalue
+$E_{\ell}(k)=4\sin^{2}\Big{(}\frac{\pi\ell}{2L}\Big{)}+4\cos\Big{(}\frac{\pi\ell}{L}\Big{)}\sin^{2}\Big{(}\frac{k}{4}\Big{)}.$ (25)
+The first term is $O(L^{-2})$, which vanishes in the thermodynamic limit. The second term is $O(k^{2})$.
+Despite the presence of gapless excitations, all the correlation functions of the form
 $\langle\Phi_{0}|\hat{\mathcal{O}}_{\bm{x}}^{\dagger}(\hat{\mathbbm{1}}-\hat{G})\hat{\mathcal{O}}^{\prime}_{\bm{y}}|\Phi_{0}\rangle$ (26)
-decays exponentially with the system size for any choice of local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}^{\prime}_{\bm{y}}$.
-Here $\hat{G}$ is the projector onto the ground states of $\hat{H}^{\mathrm{pbc}}$.
-This is because both of the two ground states of $\hat{H}^{\mathrm{pbc}}$ (the first two states in Eq.
+decay exponentially with the system size for any choice of local operators $\hat{\mathcal{O}}_{\bm{x}}$ and $\hat{\mathcal{O}}^{\prime}_{\bm{y}}$.
+Here $\hat{G}$ is the projector onto the ground states of $\hat{H}$.
+This is because both of the two ground states of $\hat{H}$ (the first two states in Eq.
 (23)) are product states.
-V.3 XYZ MG model
-As our last example, let us discuss variations of the MG model. The local Hamiltonian of the original model is given by
-$\hat{Q}_{x,x+1,x+2}$
-$\coloneqq\frac{1}{2}\hat{\mathbbm{1}}+\frac{2}{3}(\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}+\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}+\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+2}).$ (76)
-Any singlet state $|s\rangle_{x,x+1}\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ belongs to the kernel of this Hamiltonian.
-$\mathrm{ker}\,\hat{Q}_{x,x+1,x+2}$
-$=\mathrm{span}\Big{\{}|010\rangle-|001\rangle,|100\rangle-|010\rangle,$
+V.1.2 2D
+In two dimensions, the critical point $J_{c}$ is known to be $(1/2)\log(1+\sqrt{2})=0.440687\cdots$.
+Above $J>J_{c}$, the $\mathbb{Z}_{2}$ symmetry is spontaneously broken and many quasi-degenerate ground states appear, whose finite size splitting decay exponentially with the system size.
+Hence one may think that this model is a counterexample to our first conjecture that states the absence of finite size splitting in gapped FF system.
+However, it is not the case because the model is known to be gapless above $J>J_{c}$ Miyashita and Takano (1985); Henkel and Pleimling (2010).
+For example, in the limit $J\to\infty$, not only all spin up state and all spin down states, there are many other ground states with straight domain walls as illustrated in Fig 2(c)  [^2] The width of each domain must be greater than one.
+The one-dimensional excitation on the surface of a domain wall [Fig 2(d)] that precisely takes the form $|\Psi_{k,\ell}\rangle$ in Eq.
+(24) is a gapless excitation with the same excitation energy $E_{\ell}(k)$ in Eq.
+(25).
+V.2 XYZ MG model
+As another nontrivial example, let us discuss a variation of the MG model Sriram Shastry and Sutherland (1981); Shastry and Sutherland (1981); Gerhardt et al.
+(1998); Xu et al.
+(2021); Palle and Benton (2021).
+The local Hamiltonian of the original model is given by
+$\hat{Q}_{x,x+1,x+2}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}+\frac{2}{3}(\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}+\hat{\bm{s}}_{x+1}\cdot\hat{\bm{s}}_{x+2}+\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+2}).$ (27)
+Any singlet state $|s\rangle_{x,x+1}\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ belongs to the kernel of $\hat{Q}_{x,x+1,x+2}$:
+$\mathrm{span}\Big{\{}|010\rangle-|001\rangle,|100\rangle-|010\rangle,$
 […]
 The extension of this model to the XYZ coupling
-$\hat{Q}_{x,x+1,x+2}^{\prime}$
-$\coloneqq J_{0}\hat{\mathbbm{1}}+\sum_{a=x,y,z}J_{a}(\hat{s}_{x}^{a}\hat{s}_{x+1}^{a}+\hat{s}_{x+1}^{a}\hat{s}_{x+2}^{a}+\hat{s}_{x}^{a}\hat{s}_{x+2}^{a})$ (78)
+$\hat{Q}_{x,x+1,x+2}^{\prime}\coloneqq\frac{1}{4}\hat{\mathbbm{1}}+\sum_{a=x,y,z}J_{a}(\hat{s}_{x}^{a}\hat{s}_{x+1}^{a}+\hat{s}_{x+1}^{a}\hat{s}_{x+2}^{a}+\hat{s}_{x}^{a}\hat{s}_{x+2}^{a})$ (29)
 is still frustration free for some ranges of $J_{x}$, $J_{y}$, $J_{z}$ Sriram Shastry and Sutherland (1981); Shastry and Sutherland (1981); Gerhardt et al.
@@
 $J_{x}J_{y}+J_{y}J_{z}+J_{z}J_{x}=0,$ (30)
-the model exhibits additional ground state degeneracy. Following Ref. Palle and Benton (2021), we set $J_{0}=\frac{1}{4}$ and
+the model exhibits additional ground state degeneracy. Following Ref. Palle and Benton, 2021, we set
 $J_{x}$ $=\frac{1}{3}-\frac{2}{3}\cos\Big{(}2\theta-\frac{2\pi}{3}\Big{)},$ (31)
@@
 The $\theta=\pm\pi/4$ cases are unitary equivalent to the model discussed in Ref.
-Saito and Hotta (2024).
+Saito and Hotta, 2024.
 The plane-wave state $|{\Psi_{k}}\rangle$ in Eq.
-(16) is not a low-energy state unless $\theta=0$; indeed, we find
-$\langle\Psi_{k}|\hat{H}^{\mathrm{pbc}}|{\Psi_{k}}\rangle=\frac{\cos^{2}\theta}{3}(1+2\cos k)^{2}+(L-3)\sin^{2}\theta.$ (86)
+(4) of the main text is not a low-energy state unless $\theta=0$; indeed, we find
+$\langle\Psi_{k}|\hat{H}|{\Psi_{k}}\rangle=\frac{\cos^{2}\theta}{3}(1+2\cos k)^{2}+(L-3)\sin^{2}\theta.$ (37)
 To construct a low-energy excitation, we assume $L$ is an odd integer greater than three. We introduce a plane-wave state of the domain wall state
-$|\tilde{\Psi}_{k}\rangle=\frac{1}{\sqrt{Ln_{k}}}\sum_{m=0}^{L-1}e^{ikm}\hat{T}^{m}|D\rangle,$ (87)
+$|\tilde{\Psi}_{k}\rangle=\frac{1}{\sqrt{Ln_{k}}}\sum_{m=0}^{L-1}e^{ikm}\hat{T}_{1}^{m}|D\rangle,$ (38)
 $|D\rangle\coloneqq|\phi\rangle_{1,2,3}|s\rangle_{4,5}|s\rangle_{6,7}\cdots|s\rangle_{L-1,L},$ (39)
 where $n_{k}\coloneqq 1-(-2)^{-\frac{L-5}{2}}\cos k$ is the normalization factor that converges to $1$ in the large $L$ limit. Using the translation symmetry of $\hat{H}$, we find
-$\langle\tilde{\Psi}_{k}|\hat{H}|\tilde{\Psi}_{k}\rangle$ $=\frac{\big{\|}\hat{Q}_{2,3,4}(|D\rangle+\hat{T}^{2}|D\rangle)\big{\|}^{2}}{n_{k}}=\frac{2\sin^{2}k}{3n_{k}}.$ (89)
+$\langle\tilde{\Psi}_{k}|\hat{H}|\tilde{\Psi}_{k}\rangle$ $=\frac{\big{\|}\hat{Q}_{2,3,4}(|D\rangle+\hat{T}_{1}^{2}|D\rangle)\big{\|}^{2}}{n_{k}}=\frac{2\sin^{2}k}{3n_{k}}.$ (40)
 This is $O(k^{2})$, which is consistent with the result of Ref.
-Palle and Benton (2021) on Kagome lattice.
+Palle and Benton, 2021 on Kagome lattice.
 Hence, our conjecture still holds in this example as well, despite the fact that $|{\Phi_{0}}\rangle$ is not a ground state and $|{\Psi_{k}}\rangle$ is not gapless.
 VI Conclusion
-In this paper, we proposed two conjectures regarding frustration-free systems.
-The first conjecture posited that no finite-size gap opens between the degenerate ground states of frustration-free systems.
+In this work, we proposed two conjectures regarding FF systems.
+The first one posited that no finite-size splitting opens between the degenerate ground states.
 While this has been implicitly assumed in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016), we pointed out the necessity of being cautious as it is actually a non-trivial assumption.
-The second conjecture stated that in a gapless frustration-free system with translational symmetry, the dispersion relation of low-energy excitations near a certain wave number $\bm{k}_{0}$ can be bounded above by a quadratic dispersion, i.e., $O\big{(}|\bm{k}-\bm{k}_{0}|^{2},L^{-2}\big{)}$.
-We proved this conjecture for the case of an $s=1/2$ spin chain with nearest-neighbor interactions by combining the result of a seminal work by Bravyi and Gosset Bravyi and Gosset (2015) and the min-max theorem.
-Furthermore, we extended these results to two- and three-dimensional models on the cubic lattice.
-However, it became clear that this method could not be directly extended to cases where three or more neighboring spins interact simultaneously.
-Extending our results to cases where the Hilbert space dimension at each site is larger than two or to models with longer-range interactions constitutes important future work.
-Finally, there is a related, more general conjecture that the finite-size gap $\epsilon=E_{\tilde{N}_{\mathrm{deg}}+1}$ for a gapless frustration-free Hamiltonian can be bounded above by $O(L^{-2})$ regardless of the presence or absence of translational symmetry or the details of the boundary conditions.
+The second conjecture stated that in a gapless FF system with translational symmetry, the dispersion relation of low-energy excitations near a certain wave number $\bm{k}_{0}$ can be bounded above by a quadratic dispersion.
+We proved this for the case of $s=1/2$ spin models with nearest-neighbor interactions defined on the hyper cubic lattice.
+However, this method could not be directly extended to models with longer-range interactions or larger local Hilbert space dimensions.
+Extending our results to these cases constitutes important future work.
+Finally, there is a related, more general conjecture that the finite-size gap $\epsilon\coloneqq E_{\tilde{N}_{\mathrm{deg}}+1}$ for gapless FF Hamiltonians can be bounded above by $O(L^{-2})$ regardless of the presence or absence of translational symmetry or the details of the boundary conditions.
 In fact, in the accompanying paper Masaoka et al.
-(2024), we prove this statement for critical frustration-free systems in which a ground-state correlation function shows a power-law behavior.
-However, such an argument is not applicable, for example, to the uncle Hamiltonian for the GHZ state since all correlation functions decays exponentially as discussed in Sec.
-V.2.
+(2024), we prove this statement for critical FF systems in which a ground-state correlation function shows a power-law behavior.
+However, such an argument is not applicable to the uncle Hamiltonian for the GHZ state since all correlation functions decays exponentially.
 Acknowledgements.
-We thank Hosho Katsura for useful discussions.
+We thank Hosho Katsura and Zijian Wang for useful discussions.
 The work of H.W.
@@
 This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S.
-Appendix A Proof of the inequality (11)
+Appendix A Proof of the min-max principle
 Let $\mathcal{M}_{j}$ be an arbitrary $j$ dimensional subspace of the entire Hilbert space $\mathcal{H}_{D}$.
@@
 In particular, if $\langle\Phi|\hat{H}|\Phi\rangle$ and $\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ are maximized by $|\Phi\rangle_{*}\in\mathcal{M}_{j}$ and $|\Phi\rangle_{*}^{\prime}\in\mathcal{M}_{j}$, respectively, we have
-$\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ $=\langle\Phi_{*}^{\prime}|\hat{H}^{\prime}|\Phi_{*}^{\prime}\rangle$
-$\geq\langle\Phi_{*}|\hat{H}^{\prime}|\Phi_{*}\rangle$
+$\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ $=\langle\Phi_{*}^{\prime}|\hat{H}^{\prime}|\Phi_{*}^{\prime}\rangle\geq\langle\Phi_{*}|\hat{H}^{\prime}|\Phi_{*}\rangle$
 $\geq\langle\Phi_{*}|\hat{H}|\Phi_{*}\rangle=\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}|\Phi\rangle.$ (44)
-This relation holds for any $\mathcal{M}_{j}\subset\mathcal{H}_{D}$. Hence, we arrive at the inequality (11).
+This relation holds for any $\mathcal{M}_{j}\subset\mathcal{H}_{D}$. Hence, we arrive at the theorem.
 Appendix B Derivation of Eq. (8)
@@
 $\hat{U}|0\rangle=e^{i\theta}(u_{1}|0\rangle+u_{2}|1\rangle),$ (45)
-$\hat{U}|1\rangle=e^{i\theta}(-u_{2}^{*}|0\rangle+u_{1}^{*}|1\rangle),$ (95)
+$\hat{U}|1\rangle=e^{i\theta}(-u_{2}^{*}|0\rangle+u_{1}^{*}|1\rangle).$ (46)
 The inverse of $\hat{U}$ reads
@@
 $\hat{U}^{\dagger}|1\rangle=e^{-i\theta}(u_{2}^{*}|0\rangle+u_{1}|1\rangle).$ (48)
-Hence
+Under this local unitary transformation, the matrix $m_{\psi}$ defined by
+$m_{\psi}\coloneqq\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}$ (49)
+is changed to
 $m_{\psi^{\prime}}$ $=\begin{pmatrix}\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|01\rangle&\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|11\rangle\\ -\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|00\rangle&-\langle\psi|\hat{U}^{\dagger}\otimes\hat{U}^{\dagger}|10\rangle\end{pmatrix}$
-$=e^{-2i\theta}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)\end{pmatrix}$
-$=e^{-2i\theta}\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|1\rangle&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|1\rangle\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|0\rangle&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|0\rangle\end{pmatrix}$
 $=e^{-2i\theta}\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}\begin{pmatrix}u_{1}^{*}&u_{2}^{*}\\ -u_{2}&u_{1}\end{pmatrix}$
 $=e^{-2i\theta}Um_{\psi}U^{\dagger}.$ (50)
-Therefore, by choosing $U=\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}$ and $\theta$ properly, one gets the form in Eq. (25), which is equivalent to the parametrization in Eq. (20).
-Appendix C The case of $\det m_{\psi}=0$.
+Therefore, by choosing $U=\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}$ and $\theta$ properly, one gets the form
+$m_{\psi}=\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}.$ (51)
+which is equivalent to the parametrization in Eq. (6) of the main text.
+Appendix C Derivation of Eq. (16)
 Here we discuss the case $\alpha=\beta=0$ in the rank 1 model.
@@
 According to Ref.
-Bravyi and Gosset (2015), excitations under OBC are gapped in this case, which implies that excitations under PBC are also gapped.
+Bravyi and Gosset, 2015, excitations under OBC are gapped in this case, which implies that excitations under PBC are also gapped.
 When $\gamma\neq 0$, orthogonal ground states under OBC are product states with a single domain-wall:
@@
 In contrast, under PBC, domain wall states violate the boundary term and the two translation-invariant product states give the ground states.
-Thus the formula in Eq.
-(32) is applicable to this case.
 On the other hand, the $\alpha=\beta=\gamma=0$ case is exceptional.